### v1

# Cascade Speculative Drafting for Even Faster LLM Inference

Abstract
Speculative decoding enhances the efficiency of large language models (LLMs) by leveraging a draft model to draft for a larger target model to review. However, drafting in speculative decoding involves slow autoregressive generation and generating tokens of different importance with the same time allocation. These two inefficiencies lead to its suboptimal performance. To address this issue, we introduce Cascade Speculative Drafting (CS. Drafting), a novel approach that employs two types of cascades. The Vertical Cascade eliminates autoregressive generation from neural models. The Horizontal Cascade constitutes efficient time allocation in drafting with its optimality supported by our theoretical analysis. Combining both cascades, our CS. Drafting algorithm has achieved up to 72 percent additional speedup over speculative decoding in our experiments while keeping the same output distribution.
Machine Learning, ICML
[…]
1 Introduction
The advent of Large Language Models (LLMs) like GPT-4 (OpenAI, 2023) has marked a significant milestone in the field of natural language processing (NLP). These models have not only excelled in various NLP tasks but also found widespread application in user-interactive settings such as chatbots and virtual assistants. This application involves an extremely high number of users of up to hundreds of millions daily. To service in real-time at this scale, a low latency system is not only cost-saving but also crucial for keeping the service running. In addition, the sheer scale of the service making even a slight improvement over the latency of LLMs can contribute greatly to the community. Consequently, optimizing LLM latency has become a crucial and impactful research focus.
Unfortunately, LLMs currently suffer from high latency especially for long-form generation as LLM can only generate one token at each run. The challenge is especially great during applications such as chatbots or virtual assistants due to the long generation of hundreds or even thousands of tokens. An emerging solution, known as speculative decoding (Leviathan et al., 2023; Chen et al., 2023), shows potential to mitigate this issue. In speculative decoding, a draft model (which is smaller and faster) generates $k$ tokens in each step(with $k$ being a hyperparameter), and these tokens are then reviewed by a target model (which is larger and slower) in parallel. The target model will accept any tokens aligned with its output and further generate one token; the process happens all in a single run while maintaining the output distribution identical to the autoregressive one by the target model itself. As a larger number of expected accepted tokens per run indicates fewer target model runs, the efficiency of speculative decoding can be attributed to two factors: the expected number of tokens accepted by the target model per run, and the autoregressive efficiency of the draft model.
Figure 1: In a single step of speculative decoding, the figure shows the probability of acceptance of draft tokens by their positions on GSM8K datasets, evaluated on FLAN-T5-small, base and large models. The draft model generates 30 tokens at each step.
Figure 2: Cascade Speculative Sampling uses LLMs of multiple sizes to generate and verify tokens. The model at the inner layer $\mathcal{M}_{d(i)}$ generates tokens that are proposed to and verified by LLMs at the next outer layer $\mathcal{M}_{d(i+1)}$. Once the verification process is completed, the Draft $\mathcal{M}_{d(i)}$ continues to generate new tokens given the accepted prefix tokens from previous layer $\mathcal{M}_{d(i-1)}$ as input.
Herein lies a trade-off: employing a larger draft model or generating more draft tokens per target run will improve the expected number of accepted tokens. However, either of these moves will increase the latency of the draft model significantly. As a result, the maximum speed improvement that the current speculative decoding algorithm can achieve is with an optimal balance between size and number of generated tokens. From previous work by Leviathan et al. (2023), it’s shown that very small models (less than 100 times the size of the target model) are optimal for most tasks experimented representing that the cost of drafting model can be the limiting factor of the speculative decoding algorithm. This observation leads to our central idea, due to the significant latency caused by the drafting model, improving drafting efficiency is the key to improving the performance of speculative decoding.
During our preliminary experiment, we found that speculative decoding with a simple non-neural language model can certainly improve the performance of a language model. This is because a simple non-neural language model is almost free of computational meaning that it’s free of the trade-off above. However, due to the limiting ability of such a simple non-neural language model, speculative decoding does not yield optimal results with a simple non-neural language model compared to a well-sized language model from the same model family of the target model. This makes a simple non-neural language model model suboptimal when a well-sized language model is available. We organize these observations into two laws of autoregressive generation of language models:
1. Without regard to generation length nor any additional hyperparameter tuning, a simple non-neural language model can always improve the generation speed of a language model by using speculative decoding over autoregressive generation.
2. When there is a well-sized smaller model(of the same architecture, distilled from the target model or trained from a similar corpus) and a task with enough expected generation length, the smaller model can be used to improve the generation speed of the target model over the simple non-neural language model.
These two laws provide us with insight into designing a new algorithm for even further speed improvement. We propose a new algorithm Cascade Speculative Drafting (CS. Drafting) a speculative execution algorithm with a paradigm that radically excludes any autoregressive generation from neural language models and exhaustively substitutes any simple non-neural language models with superior neural language models. The resulting algorithm contains multiple draft models with the smallest being a simple non-neural language model, with each neural draft model reviewing generations from a smaller model while proposing its reviewed content to a larger draft model or the target model. We name this layered speculative decoding design the Vertical Cascade.
Moreover, our experimental findings reveal an intriguing pattern: in a drafting step, the tokens generated later in the sequence by the draft model exhibit a progressively lower probability of being accepted by the target model, as illustrated in Figure 1. (Note that this is different from the acceptance rate which is the probability of the token being accepted conditioned on the previous token being accepted) The observation suggests that the later tokens generated by draft models, with a higher likelihood of rejection, contribute much less to the expected number of accepted tokens while creating the same latency. To address this, we propose using smaller, faster draft models $\mathcal{M}_{d(i-1)}$ for generating these high-rejection tokens, forming the Horizontal Cascade. Along with the aforementioned Vertical Cascade, these strategies compose our complete CS. Drafting approach illustrated in Figure 2. Integrating these vertical and horizontal cascades, our method innovatively mitigates LLM latency issues.
We evaluated our proposed CS. Drafting on a variety of tasks and model sizes. The experimental results demonstrate the superiority of our method. Moreover, theoretical analysis is provided to prove the optimality of our proposed method. Ablation studies further verify the effectiveness of our method.
In summary, our main contributions are as follows:
  • We introduce a novel algorithm, Cascade Speculative Drafting (CS. Drafting), a speculative-execution-based algorithm improving language models inference speed without compromising on generality or overall efficiency.
  • Our work includes a theoretical analysis affirming the optimality of our CS. Drafting approach.
  • We have established a new evaluation method, standardized walltime improvement, tailored for algorithms based on speculative decoding to ensure reproducibility and facilitate consistent comparisons across different studies.
  • Through our empirical study, we demonstrate that our CS. Drafting algorithm consistently achieves lower latency compared to vanilla speculative decoding across different tasks and settings. With a speedup of up to 72% over speculative decoding, this empirical evidence underscores the practical benefits and efficiency gains of our proposed approach.
2 Preliminary
2.1 Speculative Decoding Algorithm
In real-world applications of large language models (LLMs), the sequential nature of autoregressive decoding emerges as a primary latency bottleneck in terms of response time. Addressing this, speculative decoding (SD) (Leviathan et al., 2023) presents an innovative solution that utilizes a compact draft model for sequential token generation with validation by a larger target model resulting in reduced latency.
The core concept of speculative decoding involves a novel algorithm that accelerates sampling from autoregressive models without altering output distributions. At its heart, there are two key observations: 1) certain generations in language modeling are simpler than others and can be predicted by more efficient models correctly, and 2) using speculative execution along with a new sampling method enables faster, exact decoding from large models. generation
Specifically, let $x$ be the input tokens at a run and $\mathcal{M}_{t}$ and $\mathcal{M}_{d}$ are the target and the draft model respectively. We interpret speculative sampling as a two-stage operation. In the proposing stage, we sample $\{x_{1},...,x_{k}\}$ from draft model $\mathcal{M}_{d}$ autoregressively. In the reviewing stage, let $x_{i}\in\{x_{1},...,x_{k}\}$ represents the token at the current position, and we accept it if $\mathcal{M}_{d}(x)[i-1]<\mathcal{M}_{t}(x)[i-1]$ ; in the event that $\mathcal{M}_{d}(x)[i-1]>\mathcal{M}_{t}(x)[i-1]$, we reject $x_{i}$ with a probability of $1-\frac{\mathcal{M}_{t}(x)[i-1]}{\mathcal{M}_{d}(x)[i-1]}$ and proceed to resample $x_{i}$ from a recalibrated distribution $\text{norm}(\max(0,\mathcal{M}_{t}(x)[i-1]-\mathcal{M}_{d}(x)[i-1]))$ an then reject any token following $x_{i}$. Such a design guarantees the output is the same as sampling autoregressively using the target model alone.
This method was empirically validated on various tasks and model sizes, demonstrating a significant acceleration in inference times (2x-3x faster) compared to standard implementations, without affecting the outputs. Importantly, it does not require task-specific training, altering model architectures, or changing training procedures, making it a practical solution for enhancing the efficiency of autoregressive models such as Transformers (Leviathan et al., 2023).
2.2 Notation and Definition
We introduce notions and definitions that are used throughout the paper. Let $\mathcal{M}_{t}$ be the target model, $\{\mathcal{M}_{i}\}$ be the draft models assisting generation with $\mathcal{M}_{i}$ being the draft model generating the $i$-th token. In the simpler case of the speculative decoding, $\mathcal{M}_{i}=q$ for any $i$. Let $x$ be the input to the model at a single run, $\mathcal{M}_{t}(x)$ and $\mathcal{M}_{i}(x)$ are then the output probability distribution with input $x$.
Expected acceptance rate $\alpha=\mathbf{E}(1-\frac{1}{2}|\mathcal{M}_{t}(x)-\mathcal{M}_{i}(x)|)$ is the probability of draft generation by $\{\mathcal{M}_{i}\}$ being accepted by target model $\mathcal{M}_{t}$ (Leviathan et al., 2023).
Cost coefficient $c_{i}$ is the ratio of time for a single run of $\mathcal{M}_{i}$ over $\mathcal{M}_{t}$.
Lenience in speculative decoding is a parameter to loosen the review process by the target model for further speedup with the tradeoff that the generation result might not be the same as the target model. Yet in a recursive speculative decoding setup, lenience in the inner speculative decoding will not affect the final output. With lenience $l\in[1,\infty)$, The acceptance condition of $x_{i}$ speculative decoding will become $\mathcal{M}_{d}(x)[i]<l*\mathcal{M}_{t}(x)[i]$; otherwise, with a probability of $1-\frac{l*\mathcal{M}_{t}(x)}{\mathcal{M}_{d}(x)}$ , we reject the $x_{i}$ and any following tokens.
Expected walltime improvement factor (EWIF) is the expected time improvement of an algorithm based on speculative decoding over the vanilla generation of the target model.
3 Cascaded Speculative Drafting
Algorithm 1 CascadeSpeculativeDrafting
  Require:draft model $\{\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{n}}\}$, target model $\mathcal{M}_{t}$, $prefix$, flag $isFirstCall$, hyperparameters $K_{nn}$, $l$
  draftList $\leftarrow$ $[\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{n}}]$
  curGen $\leftarrow$ $prefix$, curProbs $\leftarrow$ []
   $\triangleright$ Unpack the current parameters from
  $[k_{1},...,k_{n-1}]\leftarrow$ first row of $K_{nn}$
   $\triangleright$ Base case when the MaG generate
  if draftList only contains one element then
     $\mathcal{M}\leftarrow$ first element of draftList
     $res\leftarrow$ $\mathcal{M}(curGen)$
      $\triangleright$ MaG generates greedily, so the logits will be 1 at the position of the generation token and 0 otherwise
     return $res.generation,res.logits$
  end if
   $\triangleright$ Begin the cascade
  for $i\leftarrow 1$ to $n$ do
     curTarget $\leftarrow$ the $i$-th item of draftList
     curDraftList $\leftarrow$ the sublist of draftList starting from index $i+1$
     curK $\leftarrow$ the submatrix of $K_{nn}$ from with left-top cornor (i + 1, j) to the right-most bottom
     while length of curGen is less than $k_{i}$ do
        $[x_{1},..,x_{r}],[p_{1},p_{2},...,p_{r}]\leftarrow$ CascadeSpeculativeDecoding(curDraftList, curTarget, curGen, curK, $l$)
        curGen $\leftarrow[x_{1},..,x_{s}]$
         $\triangleright$ Do not overwrite the previous probabilities
        s = curProbs.lengths + 1
        Add elements of $[p_{s},p_{2},...,p_{r}]$ to curProbs
     end while
  end for
   $\triangleright$ No leniency when the target model reviews
  if isFirstCall then
     $l=1$
  end if
   $\triangleright$ Review process by $\mathcal{M}_{t}$ to ensure the output distribution is the same as autoregressive generation
  $[p_{1}^{\prime},p_{2}^{\prime},...,p_{m}^{\prime}]\leftarrow$ $\mathcal{M}_{t}(curGen)$
  $index\leftarrow-1$
  for $i\leftarrow 1+prefix.length$ to $m-1$ do
     Sample $r$ from a uniform distribution
     if $r\leq min(1,\frac{lp_{i}^{\prime}}{p_{i}})$ then
        $index\leftarrow i$ and exit the loop
     end if
  end for
  if index is -1 then
     $index\leftarrow m$
  end if
  Sample $x_{last}$ from $p_{index}^{\prime}-p_{index}$
  return $prefix+[x_{1},...,x_{index-1}]+[x_{last}]$, $[p_{1}^{\prime},p_{2}^{\prime},...,p_{index}^{\prime}]$
3.1 Horizontal Cascade
One key observation that leads to our experiment is that during the drafting steps of speculative decoding, not all drafting tokens are made equal, as the first draft token is exponentially easier to be accepted than the $k$-th draft token since the expected acceptance rate of the $q_{i}$ decreases exponentially as $i$ increases. In Corollary 4.3, we proved the exponential decrease above and also observed that the derivative of EWIF with respect to $\alpha_{1}$ is also exponentially greater than the derivative of EWIF with respect to $\alpha_{k}$ representing the effort in improving model acceptance rate on the first token, though having the same cost, is exponentially more valuable than the effort on the last token. As a result, given multiple models available for drafting, we propose cascade speculative decoding which involves using a larger model to generate the first few tokens and a smaller model to generate the remaining tokens.
In addition, assuming $\mathcal{M}_{d1},\mathcal{M}_{d2}$ being two draft models [^1] The notation of $\mathcal{M}_{d1}$ is different from $\mathcal{M}_{i}$ When the $d$ appeared in the subscript, it means the model can be used for generating multiple tokens. with $\mathcal{M}_{d1}$ having high cost and acceptance rate then $\mathcal{M}_{d2}$ during drafting of $i$-th token with cascade decoding, a more general setup can be achieved by using an interpolated draft model by assigning a probability $f(i)$ such that with probability $f(i)$ the $i$-th token is generated by $\mathcal{M}_{d1}$; otherwise, it’s generated by $\mathcal{M}_{d2}$. This interpolated draft model setup can provide a smoother transition of cost and acceptance rate. However, we prove that such a setup will not improve the walltime improvement factor in theorem 4.4.
3.2 Vertical Cascade
With our proven results of theorem 4.4 that for optimal cascade sampling of two draft model $\mathcal{M}_{d1},\mathcal{M}_{d2}$, each draft token will be generated by one of the draft models, we observe one additional inefficiency even with the already improved efficiency with horizontal cascade. As we have two draft models $\mathcal{M}_{d1},\mathcal{M}_{d2}$ with $\mathcal{M}_{d1}$ having better acceptance rate and higher time cost, we hypothesize that another round of speculative sampling replacing autoregressive generation of $\mathcal{M}_{d1}$ for drafting can further improve the efficiency of cascade sampling. Furthermore, leniency introduced by Leviathan et al. (Leviathan et al., 2023) provides a tradeoff between quality and speed. In practice, such a quality drop can be undesirable; however, the Cascade Speculative Drafting algorithm does not have such a limitation. Since the final quality is always guaranteed by the target model, any lenience in speculative decoding between the draft models will not affect the final output. In addition, we found that lenience can potentially bring further speedup as the optimal matching to the larger draft model in speculative drafting might not reduce the generation speed as much as the speedup introduced by lenience.
3.3 Max-gram for Better simple non-neural language model
A simple non-neural language model is special to speculative decoding as it almost incurs no cost when drafting for its target model. This means despite a slight implementation overhead, it can almost always replace autoregressive generation using greedy decoding, sampling, or beam search. In addition, in the CS Drafting algorithm, a simple non-neural language model also generates for the draft model that reviews the most tokens making the output alignment between the simple non-neural language model and the smallest draft model very important for the speedup of our algorithm.
During our pursuing of a better simple non-neural language model, we noticed a general pattern. When language models generate, some words and phrases in the input query appear repeatedly in the generation result. A common sense that the answer should have words and phrases overlapping with the question.
Based on the observation, we designed the Max-Gram(MaG) algorithm that greedily finds maximal matches in the initial input(or existing generation) and tokens from the end of the generation, and in case there is no match, we use a bigram model pretrained on the Wikipedia. (We chose Wikipedia to not affect the generality of our experiment) To our surprise, despite its simplicity, we found this simple and general algorithm very effective in aligning with language model model output. We included a GPU-friendly version of the Max-Gram algorithm in the appendix A.
3.4 Algorithm
Combining the horizontal and vertical cascades, the algorithm is shown in Algorithm 1. The center of it has the a for loop which performs the horizontal cascade while the recursive call forms the vertical cascade. Noticeably, the algorithm with the last element of the draft models being the MaG algorithm, there will be no neural language models generate autoregressively.
4 Theoretical Analysis
We perform theoretical analysis to provide simple heuristics for evaluating a CS. Drafting system, and to support the generality of CS. Drafting, we prove the optimality of horizontal cascade.
4.1 Expected Walltime Improvement
We provide the formula for EWIF in an extended setting when multiple draft models are involved.
Theorem 4.1.
the expected walltime improvement factor(EWIF) by Algorithm 1 is $T(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})=\frac{\sum_{i=0}^{k}\prod_{j=1}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$
Proof.
Since every token will only be considered by target model if all tokens before are accepted, a simple recursion leads to the probability of acceptance of the $i$-th token being $\prod_{j=1}^{i-1}\alpha_{j}$, and the $\prod_{j=1}^{0}\alpha_{j}=1$ representing the token generated by the target model. In the formula of the EWIF, the numerator is the expected number of tokens generated. The denominator is the time costs for each step of such a system. The formula is then the EWIF. ∎
We can immediately achieve a formula for speculative decoding by requiring all the draft models to be the same.
Corollary 4.2.
When $q_{1}=q_{2}=...=q_{k}$, let $\alpha_{i}=\alpha$ and $c_{i}=c$ for any ${}_{1}$, the EWIF is $\frac{1-\alpha^{k+1}}{(1-\alpha)(ck+1)}$
Furthermore, theorem 4.1 can be used to analyze the importance of the tokens in the drafting step.
Corollary 4.3.
The probability of $i$-th token being accepted is $\prod_{j=1}^{i}\alpha_{j}$. The derivative of EWIF with respect to $\alpha_{l}$ is $\frac{dT(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})}{d\alpha_{l}}=\frac{\sum_{i=l}^{k}\prod_{j=1,j\neq l}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$. Specifically, $\frac{dT(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})}{d\alpha_{1}}=\frac{\sum_{i=1}^{k}\prod_{j=2}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$ and $\frac{dT(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})}{d\alpha_{k}}=\frac{\prod_{j=1}^{k-1}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$
Using the information provided by (Leviathan et al., 2023), we calculated a simulated EWIF with the assumption that the event of acceptance by the target model is a Bernoulli trial. Results are shown in table 1 which shows that the cascade speculative sampling achieved better EWIF over vanilla speculative under the assumption.
4.2 Generality of Horizontal Cascade
The setting of cascade drafting can be simply extended using model interpolation that involves choosing between two models using randomness. Given two draft models $d_{1}$ and $d_{2}$ and a probability function $f$, $q_{i}$ is considered to be $f{i}d_{1}+(1-f(i))d_{2}$i.e with probability $f(i)$, $q_{i}$ is $d_{1}$, with probability $1-f{i}$ $q_{i}$ is $d_{2}$, as a result, $\alpha_{i}=f(i)\alpha_{d1}+(1-f(i))\alpha_{d2}=\alpha_{d2}+f(i)(\alpha_{d1}-\alpha_{d2})$ and $c_{i}=f(i)c_{d1}+(1-f(i))c_{d2}=c_{2}+f(i)(c_{d1}-c_{d2})$. In theorem 4.4, we prove that the optimal can always be reached without interpolation.
Theorem 4.4.
Given an cascade drafting system with the setting mentioned above. There exists a probability function $f$ that optimizes the system such that $f$ satisfies $f(i)\in\{0,1\}$ for any $i$. This means each of the tokens will be generated by either $d_{1}$ or $d_{2}$, and the optimal system is deterministic.
Proof of Theorem 4.4 is included in appendix B
Dataset $\mathcal{M}_{d1}$ $\mathcal{M}_{d2}$ $k$ $k_{2}$ EWIF
CNNDM T5-small - 9 - 2.65
CNNDM T5-base - 8 - 2.96
CNNDM T5-base T5-small 5 3 3.03
ENDE T5-small - 12 - 3.61
ENDE T5-base - 11 - 3.75
ENDE T5-base T5-small 5 8 3.93
Table 1: Simulated EWIF under the assumption that the acceptance distribution is a Bernoulli distribution. In the simulation, cascade drafting exceeded the performance of the speculative decoding on both CNN Dailymail and WMT EnDe datasets.
5 Evaluation
Previous works on speculative decoding and related methods relied on walltime as an evaluation method. However, there are standardization and legitimacy concerns related to walltime. Leviathan et al. (2023) mentioned that cost coefficiency, the ratio between the time for a single run of draft model and target model will vary from system to system which can result in significant variation in wall time improvement across different systems, making comparisons across different research work difficult.
In addition, a recent analysis suggests GPU speed can vary for the same GPU model with one being 1.5x faster than another GPU of the same model (Sinha et al., 2022), and even for the same GPU, the performance variation between different runs is unknown; therefore, theoretical walltime improvement up to 1.5x can be undermined by GPU variation, and methods without any improvement can be measured with an improvement up to 1.5x by running on a different GPU. The properties of such lack of standardization and stability make walltime unfit for benchmarking algorithmic improvement. To resolve this issue, we designed a new benchmarking method to measure speculative decoding and its related algorithms which can reduce the effect of different/unstable GPU performance.
Our proposed method, standardized walltime improvement(SWI) calculates the GPU times of the models, assuming that each run of a language model costs the same amount of time, an assumption made when inventing the speculative sampling algorithm (Chen et al., 2023). Given a speculative decoding system $S$ containing k draft models $\mathcal{M}_{d1},...,\mathcal{M}_{dk}$, target model $\mathcal{M}_{t}$ solving a task $T$, and let $count(S,m,T)$ be the count of runs of model $m$ during the execution of task $T$ with system $S$, let $c_{i}$ be the expected cost coefficient being the time ratio between a single run of $\mathcal{M}_{i}$ and $\mathcal{M}_{t}$, and let $n$ be the total of tokens generated, the standardized walltime improvement is defined as
$\begin{split}SWI&(\mathcal{M}_{d1},...,\mathcal{M}_{dk},\mathcal{M}_{t},T)=\\ &\frac{n}{count(S,\mathcal{M}_{t},T)+\sum_{i=1}^{n}count(S,\mathcal{M}_{di},T)c_{i}}\end{split}$
which is simply the sum of GPU time by all models when $c_{i}$ is measured during a local experiment. The benefit of the $SWI$ evaluation is that $count(S,\mathcal{M}_{t},T)$ and $count(S,\mathcal{M}_{di},T)$ will be the same across different systems, so by using the same $c_{i}$ across different experiments the GPU performance variation will be eliminated; therefore, it provides a standardized, reproducible benchmark.
During our experiment, we will use two sets of $\{c_{i}\}$, existing $\{c_{i}\}$ reported by (Leviathan et al., 2023) and an heuristic $\{c_{i}\}$ being the ratio of number of tunable parameters between $\mathcal{M}_{di}$ and $\mathcal{M}_{t}$.
Dataset Algorithm $\{\mathcal{M}_{di}\}$ SWI (MS) $k_{11}$ $k_{21}$ $k_{12}$ $l$
GSM8K Autoregressive - 1 - - - -
GSM8K S. Decoding base $3.38$ 10 - - -
GSM8K S. Decoding small $3.06$ 11 - - -
GSM8K C.S. Drafting base, mag $3.62$ 9 - - -
GSM8K C.S. Drafting small, mag $3.16$ 11 - - -
GSM8K C.S. Drafting base, small, mag 3.75 7 1 3 1.5
MMLU Autoregressive - 1 - - - -
MMLU S. Decoding base $3.97$ 13 - - -
MMLU S. Decoding small $4.12$ 19 - - -
MMLU C.S. Drafting base, mag $4.50$ 12 - - -
MMLU C.S. Drafting small, mag $4.36$ 17 - - -
MMLU C.S. Drafting base, small, mag 4.83 5 1 14 2
Dataset Algorithm $\{\mathcal{M}_{di}\}$ SWI (PW) $k_{11}$ $k_{21}$ $k_{12}$ $l$
GSM8K Autoregressive - 1 - - - -
GSM8K S. Decoding base $2.99$ 8 - - -
GSM8K S. Decoding small $2.76$ 8 - - -
GSM8K C.S. Drafting base, mag $3.22$ 8 - - -
GSM8K C.S. Drafting small, mag $2.90$ 7 - - -
GSM8K C.S. Drafting base, small, mag 3.29 5 1 4 1.5
MMLU Autoregressive - 1 - - - -
MMLU S. Decoding base $3.42$ 10 - - -
MMLU S. Decoding small $3.51$ 11 - - -
MMLU C.S. Drafting base, mag $4.07$ 9 - - -
MMLU C.S. Drafting small, mag $3.97$ 12 - - -
MMLU C.S. Drafting base, small, mag 4.23 5 1 8 1.5
Table 2: The experiment results from dataset GSM8K and MMLU. BASE and SMALL refer to FLAN-T5-base and FLAN-T5-small. SWI(model size) is the standardized walltime improvement with the assumption the latency of each run of a model is its number of parameters(MS) and time cost data from previous work(PW)(Leviathan et al., 2023) respectively.
6 Experiment
6.1 Setup
We choose our target and draft models from FLAN-T5 (Chung et al., 2022) family for our experiment as there are large variations in model sizes in the FLAN-T5 family (from 77 million to 11 billion); also, the reasoning ability of the FLAN-T5 models providing more options for our experiments. To make our experiment similar to real-world usage, We do not perform fine-tuning and the generation is done in a zero-shot manner. We use FLAN-T5-xxl as our target model, and T5-base and T5-small as our reviewing draft models and Max-Gram algorithm to be the generating draft model.
With limited computational budge, we are only able to perform sampling in temperate 00 (greedy decoding) due to the extra steps of hyperparameter tuning needed for the Cascade Speculative Decoding algorithm; the reason behind this is the benefit of temperate 00 that we can reuse the target model output which is deterministic which saves a great amount of computational resource on running the target models repeatedly.
Cascade Speculative Drafting Inference To simplify the number of hyperparameters to tune, we use MaG to generate 10 tokens at once as it’s rare that more than 10 tokens will be accepted, and we do not use lenience when the reviewer is $\mathcal{M}_{t}$ to ensure the output distribution remains do not change, and we also eliminate lenience between $\mathcal{M}_{d2}$ and MaG due to the performance map between them are large.
With these constraints, we are left with at most four parameters $k_{11}$, $k_{21}$, $k_{12}$, $l$. For the step of CS Drafting when the target model is the reviewer $k_{11}$ and $k_{12}$ are used. $k_{21}$, $l$ in the step when $\mathcal{M}_{d1}$ is the reviewer. However, there are still too many hyperparameters for us to search for the optimal result exhaustively; instead, we use halving grid search to search for a set of hyperparameters to compare to the baseline.
Baseline We consider speculative decoding and vanilla autoregressive generation as our baseline. For speculative decoding, we consider T5-small and T5-base as the draft model as they perform better than T5-large in previous work(Leviathan et al., 2023) and our preliminary experiment. We try a large range of k-values from 2 to 30 to ensure that we will find the best result for each model as the draft model to serve as our baseline.
Dataset We chose reasoning tasks for our experiment as they present more challenges of real-world NLP applications than summarization tasks. For both datasets, we experiment in a zero-shot chain-of-thought (Wei et al., 2023) setup.
The datasets we experimented on are the following:
GSM8k (Cobbe et al., 2021) A challenging mathematical reasoning dataset many large language models struggle with.
MMLU (Hendrycks et al., 2021) A reasoning dataset in many different subjects. MMLU does not have training sets meaning applying task-specific methods such as knowledge distillation to improve speculative decoding can be challenging.
6.2 Experimental Results
The experiments are shown in table 2. In two settings of SWI and Walltime, Cascade Speculative Decoding has exceeded the speculative decoding algorithm. For GSM8K, Cascade Speculative Decoding has a maximum improvement of $37\%$ over the fastest speculative algorithm; for MMLU, the maximum improvement of Cascade Speculative Decoding over speculative decoding is $72\%$.
One other observation is that despite during speculative decoding t5-small has out-performed t5-base on some experiments, but in CS Drafting with only one neural model and MaG, T5-base consistently outperforms t5-small.
7 Related Work
7.1 Efficienct Methods for Natural Language Inference
In the era of large language models, the efficiency during inference becomes a key to model service. To reduce the model inference cost and speed up, several efficient methods have been proposed, including pruning, knowledge distillation and quantization (Treviso et al., 2023). Model pruning takes structured (Xia et al., 2022; Voita et al., 2019) or unstructured (Guo et al., 2021; Chen et al., 2020) methods to remove the redundant model parameters to reduce the storage memory and increase inference speed. Knowledge distillation takes the approach of transferring knowledge from a superior teacher model to a smaller student model (Hinton et al., 2015; Gou et al., 2021). Quantization maps high-precision data representations (e.g. 32 bits) into low-precision ones (e.g. 8 bits) to reduce memory consumption (Bhandare et al., 2019; Schaefer et al., 2023).
7.2 Speculative Decoding
Recent works improved the speed of speculative decoding by experimenting with task-specific knowledge distillation (Zhou et al., 2023), or using approaches similar to speculative decoding to achieve a tradeoff between quality and speed (Leviathan et al., 2023; Kim et al., 2023). However, the reduced quality for speed can be undesirable, and task-specific training might not be applicable to many real-world NLP applications such as chatbots; overall, we found no generic improvement for speculative decoding in the current literature.
Speculative decoding (Chen et al., 2023; Leviathan et al., 2023) is proposed to improve inference speed by using a draft model to perform auto-regressive decoding and a target model to check and replace the generated tokens. A key to further speed up is to increase the acceptance rate in speculative decoding, which can be done by aligning the distribution between the draft model and the target model. Some recent work has proposed to improve the performance of speculative decoding in a domain/task-specific setting. Zhou et al. (2023) use generalized knowledge distillation to finetune the draft model and achieve 10%-45% speedup compared to draft models without knowledge distillation. Liu et al. (2023) propose Online Speculative Decoding to continually update multiple draft models on observed user query data, and the draft models align with the query distribution in real-time. However, both approaches require additional task-specific training making it difficult to apply the methods in a general setting. On the other hand, He et al. (2023) proposed constructing a datastore from a domain-specific corpus to serve as a draft model which can be used when a draft model is not available.
8 Conclusion
In our work, we proposed a novel algorithm CS. Drafting that involves two cascades, the vertical cascade, and the horizontal cascade. The vertical cascade eliminates the necessity of autoregressive generation from a neural language model, and the horizontal cascade effectively allocates the cost for drafting tokens at different positions. Our theories proved the optimality of the horizontal cascade, and empirically, we showed that CS. Drafting achieved up to 72 percent more speedup over speculative decoding without any additional training while maintaining generation quality.
Appendix A Max-Gram Implementation
Listing 1: Max-Gram Algorithm
1def torch_index(t, value):
2    return (t == value).nonzero(as_tuple=True)[0][0]
3
4def max_gram(input_ids, encoder_ids, n=1):
5    matches = (encoder_ids[0] == input_ids[0, -1]).int()
6    if matches.sum() < 1:
7        return None
8    for i in range(2, input_ids.shape[-1] + 1):
9        new_matches = (encoder_ids[0, :(-1 * (i - 1))] == input_ids[0, -1 * i]).int()
10        combined_matches = (2 - new_matches == matches[1:]).int()
11        if combined_matches.sum() < 1:
12            index = torch_index(torch.cat(
13                (
14                    torch.tensor([0] * (i - 1), device=torch.device(encoder_ids.device)),
15                    matches
16                ),
17                dim=-1
18            ), 1)
19            return encoder_ids[:, index:index + n]
20        else:
21            matches = combined_matches
22    index = torch_index(torch.cat((
23        torch.tensor([0] * (encoder_ids.shape[-1] - matches.shape[-1])), matches), dim=-1
24        ), 1)
25    return encoder_ids[:, index+1:index + n+1]
Appendix B Proof and Analysis of Theorem 4.4
We first prove a lemma that helps the proof of theorem 4.4
Lemma B.1.
$\forall a,b,c,d\in\mathbb{R}_{>0},\gamma\in(0,1)$
$\frac{\gamma a+(1-\gamma)b}{\gamma c+(1-\gamma)d}\leq\max(\frac{a}{c},\frac{b}{d})$
Proof.
Without loss of generality, assume $\frac{a}{c}\geq\frac{b}{d}$,
$\frac{a}{c}\geq\frac{\gamma a+(1-\gamma)b}{\gamma c+(1-\gamma)d}$
$\Longleftrightarrow a(\gamma c+(1-\gamma)d)\geq c(\gamma a+(1-\gamma)b)$
$\Longleftrightarrow(1-\gamma)ad\geq(1-\gamma)cb$
$\Longleftrightarrow\frac{a}{c}\geq\frac{b}{d}$
Proof of Theorem 4.4
Proof.
Let $\alpha_{d_{1}},\alpha_{d2}$ and $c_{d_{1}},c_{d2}$ be the acceptence rate and cost ratio for model $d_{1},d_{2}$ respectively. Without loss of generality, we assume $\alpha_{d_{1}}>\alpha_{d2}$ and $c_{d_{1}}>c_{d2}$ Then the system satisfies $\alpha_{i}=f(i)\alpha_{d_{1}}+(1-f(i))\alpha_{d_{2}}$ and $c_{i}=f(i)\alpha_{d_{1}}+(1-f(i))c_{d_{2}}$ We prove using contradiction. Assuming all optimal systems need at least one interpolation of models at some positions. Let an optimal system of $f$ given $d_{1},d_{2}$, Without loss of generality, assume $r$ to be the first indices such that $f(r)$ not being an integer and the system $f$ to be the one with largest $r$ among all optimal systems.
$T(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})=\frac{\alpha_{r}(1+\sum_{i=r+1}^{k}\prod_{j=r+1}^{i}\alpha_{j})+\sum_{i=0}^{r-1}\prod_{j=1}^{i}\alpha_{j}}{1+c_{r}+\sum_{i=1,i\neq r}^{k}c_{i}}$
$=\frac{(f(r)\alpha_{d_{1}}+(1-f(r))\alpha_{d_{2}})(1+\sum_{i=r+1}^{k}\prod_{j=r+1}^{i}\alpha_{j})+\sum_{i=0}^{r-1}\prod_{j=1}^{i}\alpha_{j}}{1+f(r)c_{d_{1}}+(1-f(r))c_{d_{2}}+\sum_{i=1,i\neq r}^{k}c_{i}}$
Let $A_{1}=1+\sum_{i=r+1}^{k}\prod_{j=r+1}^{i}\alpha_{j},A_{2}=\sum_{i=0}^{r-1}\prod_{j=1}^{i}\alpha_{j}$
$=\frac{f(r)(\alpha_{d_{1}}A_{1}+A_{2})+(1-f(r))(\alpha_{d_{2}}A_{1}+A_{2})}{f(r)(1+c_{d_{1}}+\sum_{i=1,i\neq r}^{k}c_{i})+(1-f(r))(1+c_{d_{2}}+\sum_{i=1,i\neq r}^{k}c_{i})}$
$\leq\max{\frac{(\alpha_{d_{1}}A_{1}+A_{2})}{1+c_{d_{1}}+\sum_{i=1,i\neq r}^{k}c_{i}},\frac{\alpha_{d_{2}}A_{1}+A_{2})}{1+c_{d_{2}}+\sum_{i=1,i\neq r}^{k}c_{i})}}$
Meaning that $f(r)=0$ or $f(r)=1$ will result in a time improvement factor no less than a non-integer $f$, meaning replacing $f(r)$ with an integer will still result in an optimal system. This contradicts with the assumption that $f$ is the system with the largest $r$. As a result, there exists an optimal system without interpolation. ∎
Figure 3: Acceptance rate of draft tokens by their positions on MMLU datasets, evaluated on FLAN-T5-small, base and large models.

### v2

@@
 Abstract
-Speculative decoding enhances the efficiency of large language models (LLMs) by leveraging a draft model to draft for a larger target model to review.
-However, drafting in speculative decoding involves slow autoregressive generation and generating tokens of different importance with the same time allocation.
+Speculative decoding reduces the latency of large language model (LLM) inference by leveraging a draft model to draft for a larger target model to review.
+However, drafting in speculative decoding involves slow autoregressive generation and allocates the same amount of time to generating tokens of varying importance.
 These two inefficiencies lead to its suboptimal performance.
-To address this issue, we introduce Cascade Speculative Drafting (CS.
-Drafting), a novel approach that employs two types of cascades.
-The Vertical Cascade eliminates autoregressive generation from neural models.
-The Horizontal Cascade constitutes efficient time allocation in drafting with its optimality supported by our theoretical analysis.
-Combining both cascades, our CS.
-Drafting algorithm has achieved up to 72 percent additional speedup over speculative decoding in our experiments while keeping the same output distribution.
+To address this issue, we introduce Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that comprises two types of cascades.
+The Vertical Cascade eliminates autoregressive generation from neural models, while the Horizontal Cascade constitutes efficient time allocation in drafting.
+Combining both cascades, our CS Drafting algorithm has achieved up to a 72 percent additional speedup over speculative decoding in our experiments, while maintaining the same output distribution as the target model.
 Machine Learning, ICML
@@
 1 Introduction
-The advent of Large Language Models (LLMs) like GPT-4 (OpenAI, 2023) has marked a significant milestone in the field of natural language processing (NLP).
-These models have not only excelled in various NLP tasks but also found widespread application in user-interactive settings such as chatbots and virtual assistants.
-This application involves an extremely high number of users of up to hundreds of millions daily.
-To service in real-time at this scale, a low latency system is not only cost-saving but also crucial for keeping the service running.
-In addition, the sheer scale of the service making even a slight improvement over the latency of LLMs can contribute greatly to the community.
+The advent of Large Language Models (LLMs), like GPT-4 (OpenAI, 2023) has marked a significant milestone in the field of natural language processing (NLP).
+These models have not only excelled in various NLP tasks but have also found widespread application in user-interactive settings such as chatbots and virtual assistants.
+This application involves an extremely high number of users, up to hundreds of millions daily.
+To serve in real-time at this scale, a low-latency system is not only cost-saving but also crucial for keeping the service running.
+In addition, the sheer scale of the service means that even a slight improvement in the latency of LLMs can greatly contribute to both the provider and the community.
 Consequently, optimizing LLM latency has become a crucial and impactful research focus.
-Unfortunately, LLMs currently suffer from high latency especially for long-form generation as LLM can only generate one token at each run.
-The challenge is especially great during applications such as chatbots or virtual assistants due to the long generation of hundreds or even thousands of tokens.
+Unfortunately, LLMs currently suffer from high latency, especially in long-form generation, as autoregressive LLMs generate tokens one by one.
 An emerging solution, known as speculative decoding (Leviathan et al., 2023; Chen et al., 2023), shows potential to mitigate this issue.
@@
 The target model will accept any tokens aligned with its output and further generate one token; the process happens all in a single run while maintaining the output distribution identical to the autoregressive one by the target model itself.
-As a larger number of expected accepted tokens per run indicates fewer target model runs, the efficiency of speculative decoding can be attributed to two factors: the expected number of tokens accepted by the target model per run, and the autoregressive efficiency of the draft model.
-Figure 1: In a single step of speculative decoding, the figure shows the probability of acceptance of draft tokens by their positions on GSM8K datasets, evaluated on FLAN-T5-small, base and large models.
+As a larger number of expected accepted tokens per run indicates fewer target model runs, the efficiency of speculative decoding can be attributed to two main factors: the expected number of tokens accepted by the target model per run, and the latency of drafting.
+Figure 1: The probability of acceptance of draft tokens in relation to their positions in a single step of speculative decoding, evaluated on FLAN-T5-small, base, and large models.
 The draft model generates 30 tokens at each step.
 Figure 2: Cascade Speculative Sampling uses LLMs of multiple sizes to generate and verify tokens.
-The model at the inner layer $\mathcal{M}_{d(i)}$ generates tokens that are proposed to and verified by LLMs at the next outer layer $\mathcal{M}_{d(i+1)}$.
-Once the verification process is completed, the Draft $\mathcal{M}_{d(i)}$ continues to generate new tokens given the accepted prefix tokens from previous layer $\mathcal{M}_{d(i-1)}$ as input.
-Herein lies a trade-off: employing a larger draft model or generating more draft tokens per target run will improve the expected number of accepted tokens.
-However, either of these moves will increase the latency of the draft model significantly.
-As a result, the maximum speed improvement that the current speculative decoding algorithm can achieve is with an optimal balance between size and number of generated tokens.
-From previous work by Leviathan et al.
-(2023), it’s shown that very small models (less than 100 times the size of the target model) are optimal for most tasks experimented representing that the cost of drafting model can be the limiting factor of the speculative decoding algorithm.
-This observation leads to our central idea, due to the significant latency caused by the drafting model, improving drafting efficiency is the key to improving the performance of speculative decoding.
-During our preliminary experiment, we found that speculative decoding with a simple non-neural language model can certainly improve the performance of a language model.
-This is because a simple non-neural language model is almost free of computational meaning that it’s free of the trade-off above.
-However, due to the limiting ability of such a simple non-neural language model, speculative decoding does not yield optimal results with a simple non-neural language model compared to a well-sized language model from the same model family of the target model.
-This makes a simple non-neural language model model suboptimal when a well-sized language model is available.
-We organize these observations into two laws of autoregressive generation of language models:
-1.
-Without regard to generation length nor any additional hyperparameter tuning, a simple non-neural language model can always improve the generation speed of a language model by using speculative decoding over autoregressive generation.
-2.
-When there is a well-sized smaller model(of the same architecture, distilled from the target model or trained from a similar corpus) and a task with enough expected generation length, the smaller model can be used to improve the generation speed of the target model over the simple non-neural language model.
-These two laws provide us with insight into designing a new algorithm for even further speed improvement.
-We propose a new algorithm Cascade Speculative Drafting (CS.
-Drafting) a speculative execution algorithm with a paradigm that radically excludes any autoregressive generation from neural language models and exhaustively substitutes any simple non-neural language models with superior neural language models.
-The resulting algorithm contains multiple draft models with the smallest being a simple non-neural language model, with each neural draft model reviewing generations from a smaller model while proposing its reviewed content to a larger draft model or the target model.
-We name this layered speculative decoding design the Vertical Cascade.
-Moreover, our experimental findings reveal an intriguing pattern: in a drafting step, the tokens generated later in the sequence by the draft model exhibit a progressively lower probability of being accepted by the target model, as illustrated in Figure 1.
-(Note that this is different from the acceptance rate which is the probability of the token being accepted conditioned on the previous token being accepted) The observation suggests that the later tokens generated by draft models, with a higher likelihood of rejection, contribute much less to the expected number of accepted tokens while creating the same latency.
-To address this, we propose using smaller, faster draft models $\mathcal{M}_{d(i-1)}$ for generating these high-rejection tokens, forming the Horizontal Cascade.
-Along with the aforementioned Vertical Cascade, these strategies compose our complete CS.
-Drafting approach illustrated in Figure 2.
-Integrating these vertical and horizontal cascades, our method innovatively mitigates LLM latency issues.
-We evaluated our proposed CS.
-Drafting on a variety of tasks and model sizes.
-The experimental results demonstrate the superiority of our method.
-Moreover, theoretical analysis is provided to prove the optimality of our proposed method.
-Ablation studies further verify the effectiveness of our method.
-In summary, our main contributions are as follows:
-  • We introduce a novel algorithm, Cascade Speculative Drafting (CS.
-Drafting), a speculative-execution-based algorithm improving language models inference speed without compromising on generality or overall efficiency.
-  • Our work includes a theoretical analysis affirming the optimality of our CS. Drafting approach.
-  • We have established a new evaluation method, standardized walltime improvement, tailored for algorithms based on speculative decoding to ensure reproducibility and facilitate consistent comparisons across different studies.
-  • Through our empirical study, we demonstrate that our CS.
-Drafting algorithm consistently achieves lower latency compared to vanilla speculative decoding across different tasks and settings.
-With a speedup of up to 72% over speculative decoding, this empirical evidence underscores the practical benefits and efficiency gains of our proposed approach.
+The model at the inner layer $\mathcal{M}_{d(i)}$ generates tokens proposed to and verified by LLMs at the next outer layer, $\mathcal{M}_{d(i+1)}$.
+This recursive process forms the vertical cascade.
+After a vertical cascade, a smaller $\mathcal{M}_{d(i)}$ will serve as the target model and perform a vertical cascade.
+The process repeats until the smaller model becomes a statistical language model that will generate autoregressively forming the horizontal cascade.
+[…]
+In light of this, there exist a type of draft models, including statistical language models like bigram language models, that are immune to drafting efficiency issues.
+Owing to their simple structure, these models incur negligible latency and computational resource costs compared to neural language models, thereby avoiding the aforementioned trade-off.
+Therefore, assuming enough parallel computing power to perform computation on additional $k$ tokens with similar time cost (Leviathan et al., 2023), utilizing speculative decoding with a statistical language model as the draft model can generally and intuitively enhance the efficiency of a neural language model.
+However, due to their limited capabilities, speculative decoding with statistical language models alone may not yield optimal results compared to using a well-sized neural language model from the same family as the target model when the neural language model is available.
+Based on the above analysis, we propose Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that comprises multiple draft models, starting with the smallest being a statistical language model.
+Each neural draft model reviews generations from a smaller model and then proposes its reviewed content to either a larger draft model or the target model.
+We refer to this tiered speculative decoding approach as the Vertical Cascade.
+Furthermore, our analysis in Figure 1 reveals an interesting pattern in the expected acceptance rate during the drafting step: tokens generated later in the sequence by the draft model show a progressively lower probability of being accepted by the target model [^1] A theoretical analysis is provided in Section 4.
+Note that this probability differs from the acceptance rate, which is the probability of a token being accepted conditioned on the previous tokens’ acceptance..
+This indicates that later tokens from draft models, more prone to rejection, contribute less to the expected number of accepted tokens, yet incur the same latency.
+Based on this observation and our practice, we suggest the use of smaller, faster draft models for generating these high-rejection tokens, forming the Horizontal Cascade.
+Along with the aforementioned Vertical Cascade, these strategies compose our complete CS Drafting approach, as illustrated in Figure 2.
+Through theoretical analysis and empirical studies, we demonstrate that the CS Drafting algorithm outperforms speculative decoding in terms of latency across various tasks and settings, achieving an additional speedup of up to 72% over speculative decoding.
+These findings highlight the practical advantages and efficiency enhancements offered by both vertical and horizontal cascades.
+Our main contributions are summarized as follows:
+  • We introduce Cascade Speculative Drafting (CS Drafting), a speculative-execution-based algorithm that improves language model inference speed without sacrificing generation quality.
+  • We provide theoretical analyses supporting the effectiveness of the proposed CS Drafting approach.
+  • We conduct empirical experiments showing that CS Drafting achieves further speedup over speculative decoding across different tasks and settings.
 2 Preliminary
@@
 At its heart, there are two key observations: 1) certain generations in language modeling are simpler than others and can be predicted by more efficient models correctly, and 2) using speculative execution along with a new sampling method enables faster, exact decoding from large models.
-generation
-Specifically, let $x$ be the input tokens at a run and $\mathcal{M}_{t}$ and $\mathcal{M}_{d}$ are the target and the draft model respectively.
+Specifically, let $x$ be the input tokens at a run and $\mathcal{M}_{t}$ and $\mathcal{M}_{d}$ are the target and the draft model respectively, and $\mathcal{M}_{t}(x)[i]$ and $\mathcal{M}_{d}(x)[i]$ be their probability output at $i$-th token when input is $x$.
 We interpret speculative sampling as a two-stage operation.
-In the proposing stage, we sample $\{x_{1},...,x_{k}\}$ from draft model $\mathcal{M}_{d}$ autoregressively.
-In the reviewing stage, let $x_{i}\in\{x_{1},...,x_{k}\}$ represents the token at the current position, and we accept it if $\mathcal{M}_{d}(x)[i-1]<\mathcal{M}_{t}(x)[i-1]$ ; in the event that $\mathcal{M}_{d}(x)[i-1]>\mathcal{M}_{t}(x)[i-1]$, we reject $x_{i}$ with a probability of $1-\frac{\mathcal{M}_{t}(x)[i-1]}{\mathcal{M}_{d}(x)[i-1]}$ and proceed to resample $x_{i}$ from a recalibrated distribution $\text{norm}(\max(0,\mathcal{M}_{t}(x)[i-1]-\mathcal{M}_{d}(x)[i-1]))$ an then reject any token following $x_{i}$.
-Such a design guarantees the output is the same as sampling autoregressively using the target model alone.
+In the proposing stage, we sample $\{x_{t+1},...,x_{t+k}\}$ from draft model $\mathcal{M}_{d}$ autoregressively and append them to $x$ In the reviewing stage, let $x_{i}\in\{x_{t+1},...,x_{t+k}\}$ represents the token at the current position, and we accept it if $\mathcal{M}_{d}(x)[i-1]<\mathcal{M}_{t}(x)[i-1]$; in the event that $\mathcal{M}_{d}(x)[i-1]>\mathcal{M}_{t}(x)[i-1]$, we reject $x_{i}$ with a probability of $1-\frac{\mathcal{M}_{t}(x)[i-1]}{\mathcal{M}_{d}(x)[i-1]}$ and proceed to resample $x_{i}$ from a recalibrated distribution $\text{norm}(\max(0,\mathcal{M}_{t}(x)[i-1]-\mathcal{M}_{d}(x)[i-1]))$ and reject any token following $x_{i}$.
+At the end, the target model will generate one additional token following the accepted tokens.
+Such a design guarantees the output is the same as sampling autoregressively using the target model alone (Leviathan et al., 2023).
 This method was empirically validated on various tasks and model sizes, demonstrating a significant acceleration in inference times (2x-3x faster) compared to standard implementations, without affecting the outputs.
-Importantly, it does not require task-specific training, altering model architectures, or changing training procedures, making it a practical solution for enhancing the efficiency of autoregressive models such as Transformers (Leviathan et al., 2023).
-2.2 Notation and Definition
-We introduce notions and definitions that are used throughout the paper.
-Let $\mathcal{M}_{t}$ be the target model, $\{\mathcal{M}_{i}\}$ be the draft models assisting generation with $\mathcal{M}_{i}$ being the draft model generating the $i$-th token.
-In the simpler case of the speculative decoding, $\mathcal{M}_{i}=q$ for any $i$.
-Let $x$ be the input to the model at a single run, $\mathcal{M}_{t}(x)$ and $\mathcal{M}_{i}(x)$ are then the output probability distribution with input $x$.
-Expected acceptance rate $\alpha=\mathbf{E}(1-\frac{1}{2}|\mathcal{M}_{t}(x)-\mathcal{M}_{i}(x)|)$ is the probability of draft generation by $\{\mathcal{M}_{i}\}$ being accepted by target model $\mathcal{M}_{t}$ (Leviathan et al., 2023).
-Cost coefficient $c_{i}$ is the ratio of time for a single run of $\mathcal{M}_{i}$ over $\mathcal{M}_{t}$.
-Lenience in speculative decoding is a parameter to loosen the review process by the target model for further speedup with the tradeoff that the generation result might not be the same as the target model.
-Yet in a recursive speculative decoding setup, lenience in the inner speculative decoding will not affect the final output.
-With lenience $l\in[1,\infty)$, The acceptance condition of $x_{i}$ speculative decoding will become $\mathcal{M}_{d}(x)[i]<l*\mathcal{M}_{t}(x)[i]$; otherwise, with a probability of $1-\frac{l*\mathcal{M}_{t}(x)}{\mathcal{M}_{d}(x)}$ , we reject the $x_{i}$ and any following tokens.
-Expected walltime improvement factor (EWIF) is the expected time improvement of an algorithm based on speculative decoding over the vanilla generation of the target model.
+Importantly, it does not require task-specific training, altering model architectures, or changing training procedures, making it a practical solution for enhancing the efficiency of LLMs.
 3 Cascaded Speculative Drafting
@@
   draftList $\leftarrow$ $[\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{n}}]$
-  curGen $\leftarrow$ $prefix$, curProbs $\leftarrow$ []
-   $\triangleright$ Unpack the current parameters from
+   $\triangleright$ Initialize curGen and curProb
+  curGen $\leftarrow$ $prefix$, curProbs $\leftarrow$ a list of ones of length prefix
+   $\triangleright$ Unpack the a list of $k$ for current function call
   $[k_{1},...,k_{n-1}]\leftarrow$ first row of $K_{nn}$
-   $\triangleright$ Base case when the MaG generate
+   $\triangleright$ Base case of the recursive call when the MaG generate
   if draftList only contains one element then
@@
      curDraftList $\leftarrow$ the sublist of draftList starting from index $i+1$
-     curK $\leftarrow$ the submatrix of $K_{nn}$ from with left-top cornor (i + 1, j) to the right-most bottom
-     while length of curGen is less than $k_{i}$ do
-        $[x_{1},..,x_{r}],[p_{1},p_{2},...,p_{r}]\leftarrow$ CascadeSpeculativeDecoding(curDraftList, curTarget, curGen, curK, $l$)
-        curGen $\leftarrow[x_{1},..,x_{s}]$
+     curK $\leftarrow$ the submatrix of $K_{nn}$ from with the top-left corner at $(i+1,j)$ extending to the bottom-right corner
+     while length of curGen - prefix.length is less than $k_{i}$ do
+        $[x_{1},..,x_{r}],[p_{1},p_{2},...,p_{r}]\leftarrow$ CascadeSpeculativeDecoding(curDraftList, curTarget, curGen, False, curK, $l$)
+        curGen $\leftarrow[x_{1},..,x_{r}]$
          $\triangleright$ Do not overwrite the previous probabilities
-        s = curProbs.lengths + 1
+[…]
         Add elements of $[p_{s},p_{2},...,p_{r}]$ to curProbs
@@
   if isFirstCall then
-     $l=1$
+     $l\leftarrow 1$
   end if
@@
   $[p_{1}^{\prime},p_{2}^{\prime},...,p_{m}^{\prime}]\leftarrow$ $\mathcal{M}_{t}(curGen)$
-  $index\leftarrow-1$
+  $index\leftarrow m$
   for $i\leftarrow 1+prefix.length$ to $m-1$ do
@@
   end for
-  if index is -1 then
-     $index\leftarrow m$
-  end if
   Sample $x_{last}$ from $p_{index}^{\prime}-p_{index}$
-  return $prefix+[x_{1},...,x_{index-1}]+[x_{last}]$, $[p_{1}^{\prime},p_{2}^{\prime},...,p_{index}^{\prime}]$
+  return $[x_{1},...,x_{index-1}]+[x_{last}]$, $[p_{1}^{\prime},p_{2}^{\prime},...,p_{index}^{\prime}]$
+3.1 Vertical Cascade
+The speculative decoding algorithm relies on the autoregressive generation of a smaller draft model.
+However, despite its smaller size, the draft model must run $k$ times for each target model run, and the cost can still be significant.
+To further reduce latency in the drafting process, we propose using an even smaller model to assist in drafting and employing the original draft model to review the generation of this smaller model.
+This process can be done recursively until it reaches the smallest draft model.
+The smallest draft model can be a statistical language model, such as a bigram language model, where the cost is negligible compared to neural language models.
+We refer to this recursive process as Vertical Cascade.
+Furthermore, to further improve speed, we can introduce leniency, a hyperparameter in speculative decoding that loosens the review process by the target model, allowing for faster speed at the trade-off of potentially differing results from the target model (Leviathan et al., 2023).
+With leniency $l\in[1,\infty)$, the acceptance condition for token $x_{i}$ in speculative decoding becomes $\mathcal{M}{d}(x)[i]<l\times\mathcal{M}{t}(x)[i]$.
+Otherwise, with a probability of $1-\frac{l\times\mathcal{M}{t}(x)}{\mathcal{M}{d}(x)}$, we reject $x_{i}$ and any following tokens.
+While the reduced quality introduced by leniency in speculative decoding is generally undesirable in practice, it fortunately does not affect the final output in our vertical cascade approach—we can only apply leniency in the inner speculative decoding.
+As long as leniency is not applied to the target model, the final quality is always guaranteed by the target model.
 3.2 Horizontal Cascade
-One key observation that leads to our experiment is that during the drafting steps of speculative decoding, not all drafting tokens are made equal, as the first draft token is exponentially easier to be accepted than the $k$-th draft token since the expected acceptance rate of the $q_{i}$ decreases exponentially as $i$ increases.
-In Corollary 4.3, we proved the exponential decrease above and also observed that the derivative of EWIF with respect to $\alpha_{1}$ is also exponentially greater than the derivative of EWIF with respect to $\alpha_{k}$ representing the effort in improving model acceptance rate on the first token, though having the same cost, is exponentially more valuable than the effort on the last token.
-As a result, given multiple models available for drafting, we propose cascade speculative decoding which involves using a larger model to generate the first few tokens and a smaller model to generate the remaining tokens.
-In addition, assuming $\mathcal{M}_{d1},\mathcal{M}_{d2}$ being two draft models [^1] The notation of $\mathcal{M}_{d1}$ is different from $\mathcal{M}_{i}$ When the $d$ appeared in the subscript, it means the model can be used for generating multiple tokens.
-with $\mathcal{M}_{d1}$ having high cost and acceptance rate then $\mathcal{M}_{d2}$ during drafting of $i$-th token with cascade decoding, a more general setup can be achieved by using an interpolated draft model by assigning a probability $f(i)$ such that with probability $f(i)$ the $i$-th token is generated by $\mathcal{M}_{d1}$; otherwise, it’s generated by $\mathcal{M}_{d2}$.
-This interpolated draft model setup can provide a smoother transition of cost and acceptance rate.
-However, we prove that such a setup will not improve the walltime improvement factor in theorem 4.4.
-3.2 Vertical Cascade
-With our proven results of theorem 4.4 that for optimal cascade sampling of two draft model $\mathcal{M}_{d1},\mathcal{M}_{d2}$, each draft token will be generated by one of the draft models, we observe one additional inefficiency even with the already improved efficiency with horizontal cascade.
-As we have two draft models $\mathcal{M}_{d1},\mathcal{M}_{d2}$ with $\mathcal{M}_{d1}$ having better acceptance rate and higher time cost, we hypothesize that another round of speculative sampling replacing autoregressive generation of $\mathcal{M}_{d1}$ for drafting can further improve the efficiency of cascade sampling.
-Furthermore, leniency introduced by Leviathan et al.
-(Leviathan et al., 2023) provides a tradeoff between quality and speed.
-In practice, such a quality drop can be undesirable; however, the Cascade Speculative Drafting algorithm does not have such a limitation.
-Since the final quality is always guaranteed by the target model, any lenience in speculative decoding between the draft models will not affect the final output.
-In addition, we found that lenience can potentially bring further speedup as the optimal matching to the larger draft model in speculative drafting might not reduce the generation speed as much as the speedup introduced by lenience.
-3.3 Max-gram for Better simple non-neural language model
-A simple non-neural language model is special to speculative decoding as it almost incurs no cost when drafting for its target model.
-This means despite a slight implementation overhead, it can almost always replace autoregressive generation using greedy decoding, sampling, or beam search.
-In addition, in the CS Drafting algorithm, a simple non-neural language model also generates for the draft model that reviews the most tokens making the output alignment between the simple non-neural language model and the smallest draft model very important for the speedup of our algorithm.
-During our pursuing of a better simple non-neural language model, we noticed a general pattern.
-When language models generate, some words and phrases in the input query appear repeatedly in the generation result.
-A common sense that the answer should have words and phrases overlapping with the question.
-Based on the observation, we designed the Max-Gram(MaG) algorithm that greedily finds maximal matches in the initial input(or existing generation) and tokens from the end of the generation, and in case there is no match, we use a bigram model pretrained on the Wikipedia.
-(We chose Wikipedia to not affect the generality of our experiment) To our surprise, despite its simplicity, we found this simple and general algorithm very effective in aligning with language model model output.
-We included a GPU-friendly version of the Max-Gram algorithm in the appendix A.
+As illustrated in Figure 1, another key observation is that during the drafting steps of speculative decoding, not all drafting tokens are created equal.
+The first draft token is more likely to be accepted as it only depends on itself; the last token is rarely accepted, as it has a chance of being reviewed only if all preceding tokens are accepted.
+From another perspective, one could also say that the first draft token bears the responsibility for all succeeding tokens, while the last token bears none.
+In Section 4, we provide a theoretical analysis of this phenomenon.
+Based on this observation, we propose better allocation of the time cost of drafting by using larger models to generate the preceding draft tokens and smaller models for the succeeding tokens.
+This design allows more time to be allocated to more effective tokens, leading to a reduction in overall latency.
+We refer to this design as the Horizontal Cascade.
+3.3 Max-gram for Better Statistical Drafting
+A statistical language model is special to speculative decoding, as it incurs almost no cost when drafting for its target model.
+In our pursuit of a more effective statistical language model, we noticed a general pattern: in language model generation, some words and phrases from the input query frequently reappear in the generated content.
+Based on this observation, we designed the Max-Gram (MaG) algorithm.
+It greedily identifies maximal matches between the initial input (or existing generation) and tokens from the end of the generation.
+In cases where there is no match, we resort to a bigram model based on the probability distribution of Wikipedia (chosen to maintain the generality of our experiment).
+Surprisingly, despite its simplicity, this straightforward and general algorithm has proven highly effective in aligning with neural language model outputs.
+We include a GPU-friendly version of the Max-Gram algorithm in Appendix A.
 3.4 Algorithm
-Combining the horizontal and vertical cascades, the algorithm is shown in Algorithm 1.
-The center of it has the a for loop which performs the horizontal cascade while the recursive call forms the vertical cascade.
-Noticeably, the algorithm with the last element of the draft models being the MaG algorithm, there will be no neural language models generate autoregressively.
+Combining the horizontal and vertical cascades, the algorithm is presented in Algorithm 1.
+At its center, there is a for-loop that performs the horizontal cascade, while the recursive call forms the vertical cascade.
+Notably, the last element of the draft models is the MaG model.
+The algorithm requires a two-dimensional hyperparameter, $K_{nn}$, with each row serving as the stop criteria for a layer of recursive calls.
+For simplicity, we assume the leniency parameter $l$ is universal for the algorithm, except when the target model is under review; thus, the algorithm can benefit from the speedup of leniency without altering the output distribution.
 4 Theoretical Analysis
-We perform theoretical analysis to provide simple heuristics for evaluating a CS. Drafting system, and to support the generality of CS. Drafting, we prove the optimality of horizontal cascade.
+In this section, we present some theoretical analyses of the proposed Cascade Speculative Drafting algorithm.
 4.1 Expected Walltime Improvement
-We provide the formula for EWIF in an extended setting when multiple draft models are involved.
+To assist our theoretical analysts, we establish the notions.
+Let $\mathcal{M}_{t}$ be the target model, $\{\mathcal{M}_{i}\}$ be the draft models assisting generation with $\mathcal{M}_{i}$ being the draft model generating the $i$-th token.
+In the simpler case of the speculative decoding, $\mathcal{M}_{i}=\mathcal{M}_{d}$ for any $i$.
+Let $x$ be the input to the model at a single run, $\mathcal{M}_{t}(x)$ and $\mathcal{M}_{i}(x)$ are then the output probability distribution with input $x$.
+Expected acceptance rate $\alpha=\mathbf{E}(1-\frac{1}{2}|\mathcal{M}_{t}(x)-\mathcal{M}_{i}(x)|)$ is the probability of draft generation by $\{\mathcal{M}_{i}\}$ being accepted by target model $\mathcal{M}_{t}$ (Leviathan et al., 2023).
+Cost coefficient $c_{i}$ is the ratio of time for a single run of draft model $\mathcal{M}_{i}$ over target model $\mathcal{M}_{t}$.
+Expected walltime improvement factor (EWIF) is the expected time improvement of an algorithm based on. We provide the formula for EWIF in an extended setting when multiple draft models are involved.
 Theorem 4.1.
-the expected walltime improvement factor(EWIF) by Algorithm 1 is $T(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})=\frac{\sum_{i=0}^{k}\prod_{j=1}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$
+The expected walltime improvement factor (EWIF) by Algorithm 1 is $T(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})=\frac{\sum_{i=0}^{k}\prod_{j=1}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$.
 Proof.
-Since every token will only be considered by target model if all tokens before are accepted, a simple recursion leads to the probability of acceptance of the $i$-th token being $\prod_{j=1}^{i-1}\alpha_{j}$, and the $\prod_{j=1}^{0}\alpha_{j}=1$ representing the token generated by the target model.
-In the formula of the EWIF, the numerator is the expected number of tokens generated.
-The denominator is the time costs for each step of such a system.
-The formula is then the EWIF.
+Since every token will only be considered by the target model if all tokens before are accepted, a simple recursion leads to the probability of acceptance of the $i$-th token being $\prod_{j=1}^{i-1}\alpha_{j}$, and the $\prod_{j=1}^{0}\alpha_{j}=1$ representing the token generated by the target model.
+In the right-hand side, the numerator is the expected number of tokens generated, and the denominator is the time costs for each step of such a system, so the right-hand side is the expected efficiency improvement.
 ∎
@@
 Corollary 4.2.
-When $q_{1}=q_{2}=...=q_{k}$, let $\alpha_{i}=\alpha$ and $c_{i}=c$ for any ${}_{1}$, the EWIF is $\frac{1-\alpha^{k+1}}{(1-\alpha)(ck+1)}$
+When $q_{1}=q_{2}=...=q_{k}$, let $\alpha_{i}=\alpha$ and $c_{i}=c$ for any $i$, the EWIF is $\frac{1-\alpha^{k+1}}{(1-\alpha)(c\cdot k+1)}$.
 Furthermore, theorem 4.1 can be used to analyze the importance of the tokens in the drafting step.
@@
 The derivative of EWIF with respect to $\alpha_{l}$ is $\frac{dT(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})}{d\alpha_{l}}=\frac{\sum_{i=l}^{k}\prod_{j=1,j\neq l}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$.
-Specifically, $\frac{dT(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})}{d\alpha_{1}}=\frac{\sum_{i=1}^{k}\prod_{j=2}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$ and $\frac{dT(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})}{d\alpha_{k}}=\frac{\prod_{j=1}^{k-1}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$
-Using the information provided by (Leviathan et al., 2023), we calculated a simulated EWIF with the assumption that the event of acceptance by the target model is a Bernoulli trial.
-Results are shown in table 1 which shows that the cascade speculative sampling achieved better EWIF over vanilla speculative under the assumption.
-4.2 Generality of Horizontal Cascade
+Specifically, $\frac{dT(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})}{d\alpha_{1}}=\frac{\sum_{i=1}^{k}\prod_{j=2}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$ and $\frac{dT(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})}{d\alpha_{k}}=\frac{\prod_{j=1}^{k-1}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$.
+Using the information provided by Leviathan et al.
+(2023), we calculat a simulated EWIF under the assumption that the event of acceptance by the target model is a Bernoulli trial.
+The results, shown in Table 1, indicate that speculative sampling with a horizontal cascade achieved better EWIF than vanilla speculative sampling under this assumption.
+4.2 Analysis of Model Interpolation
 The setting of cascade drafting can be simply extended using model interpolation that involves choosing between two models using randomness.
-Given two draft models $d_{1}$ and $d_{2}$ and a probability function $f$, $q_{i}$ is considered to be $f{i}d_{1}+(1-f(i))d_{2}$i.e with probability $f(i)$, $q_{i}$ is $d_{1}$, with probability $1-f{i}$ $q_{i}$ is $d_{2}$, as a result, $\alpha_{i}=f(i)\alpha_{d1}+(1-f(i))\alpha_{d2}=\alpha_{d2}+f(i)(\alpha_{d1}-\alpha_{d2})$ and $c_{i}=f(i)c_{d1}+(1-f(i))c_{d2}=c_{2}+f(i)(c_{d1}-c_{d2})$.
-In theorem 4.4, we prove that the optimal can always be reached without interpolation.
+Given two draft systems $d_{1}$ and $d_{2}$ (They can be two models generating autoregressive or two speculative decoding systems as used in vertical cascade) and a probability function $f$, an interpolation $d_{inter}$ is defined as $f(i)d_{1}+(1-f(i))d_{2}$.
+In other words, with probability $f(i)$, $d_{inter}$ is $d_{1}$, and with probability $1-f(i)$, $d_{inter}$ is $d_{2}$.
+As a result, $\alpha_{i}=f(i)\alpha_{d1}+(1-f(i))\alpha_{d2}=\alpha_{d2}+f(i)(\alpha_{d1}-\alpha_{d2})$ and $c_{i}=f(i)c_{d1}+(1-f(i))c_{d2}=c_{2}+f(i)(c_{d1}-c_{d2})$.
+An interpolation system can provide a smooth transition between two systems; however, in theorem 4.4, we prove that the optimal can always be reached without interpolation.
 Theorem 4.4.
-Given an cascade drafting system with the setting mentioned above.
+Given a horizontal cascade system with the setting mentioned above.
 There exists a probability function $f$ that optimizes the system such that $f$ satisfies $f(i)\in\{0,1\}$ for any $i$.
 This means each of the tokens will be generated by either $d_{1}$ or $d_{2}$, and the optimal system is deterministic.
-Proof of Theorem 4.4 is included in appendix B
-Dataset $\mathcal{M}_{d1}$ $\mathcal{M}_{d2}$ $k$ $k_{2}$ EWIF
-CNNDM T5-small - 9 - 2.65
-CNNDM T5-base - 8 - 2.96
-CNNDM T5-base T5-small 5 3 3.03
-ENDE T5-small - 12 - 3.61
-ENDE T5-base - 11 - 3.75
-ENDE T5-base T5-small 5 8 3.93
+Proof of Theorem 4.4 is included in appendix B.
+Dataset $\mathcal{M}_{d1}$ $\mathcal{M}_{d2}$ $k_{1}$ $k_{2}$ EWIF
+CNNDM small - 9 - 2.65
+CNNDM base - 8 - 2.96
+CNNDM base small 5 3 3.03
+ENDE small - 12 - 3.61
+ENDE base - 11 - 3.75
+ENDE base small 5 8 3.93
 Table 1: Simulated EWIF under the assumption that the acceptance distribution is a Bernoulli distribution.
-In the simulation, cascade drafting exceeded the performance of the speculative decoding on both CNN Dailymail and WMT EnDe datasets.
+BASE and SMALL refer to FLAN-T5-base and FLAN-T5-small.
+In the simulation, speculative sampling with horizontal cascade exceeded the performance of the vanilla speculative decoding on both CNN Dailymail (Nallapati et al., 2016) and WMT EnDe (Bojar et al., 2018) datasets.
 5 Evaluation
@@
 In addition, a recent analysis suggests GPU speed can vary for the same GPU model with one being 1.5x faster than another GPU of the same model (Sinha et al., 2022), and even for the same GPU, the performance variation between different runs is unknown; therefore, theoretical walltime improvement up to 1.5x can be undermined by GPU variation, and methods without any improvement can be measured with an improvement up to 1.5x by running on a different GPU.
-The properties of such lack of standardization and stability make walltime unfit for benchmarking algorithmic improvement.
-To resolve this issue, we designed a new benchmarking method to measure speculative decoding and its related algorithms which can reduce the effect of different/unstable GPU performance.
-Our proposed method, standardized walltime improvement(SWI) calculates the GPU times of the models, assuming that each run of a language model costs the same amount of time, an assumption made when inventing the speculative sampling algorithm (Chen et al., 2023).
+The properties of such lack of standardization and stability make walltime unfit for measuring algorithmic improvement.
+To resolve this issue, we design a new metric to measure speculative decoding and its related algorithms which can reduce the impact of varying/unstable GPU performance.
+Our proposed method, standardized walltime improvement (SWI), calculates the GPU times of the models, assuming that each run of a language model costs the same amount of time, an assumption made when inventing the speculative sampling algorithm (Chen et al., 2023).
 Given a speculative decoding system $S$ containing $k$ draft models $\mathcal{M}_{d1},...,\mathcal{M}_{dk}$, target model $\mathcal{M}_{t}$ solving a task $T$, and let $count(S,m,T)$ be the count of runs of model $m$ during the execution of task $T$ with system $S$, let $c_{i}$ be the expected cost coefficient being the time ratio between a single run of $\mathcal{M}_{i}$ and $\mathcal{M}_{t}$, and let $n$ be the total of tokens generated, the standardized walltime improvement is defined as
-$\begin{split}SWI&(\mathcal{M}_{d1},...,\mathcal{M}_{dk},\mathcal{M}_{t},T)=\\ &\frac{n}{count(S,\mathcal{M}_{t},T)+\sum_{i=1}^{n}count(S,\mathcal{M}_{di},T)c_{i}}\end{split}$
+$\begin{split}SWI&(\mathcal{M}_{d1},...,\mathcal{M}_{dk},\mathcal{M}_{t},T)=\\ &\frac{n}{count(S,\mathcal{M}_{t},T)+\sum_{i=1}^{n}count(S,\mathcal{M}_{di},T)c_{i}},\end{split}$
 which is simply the sum of GPU time by all models when $c_{i}$ is measured during a local experiment.
-The benefit of the $SWI$ evaluation is that $count(S,\mathcal{M}_{t},T)$ and $count(S,\mathcal{M}_{di},T)$ will be the same across different systems, so by using the same $c_{i}$ across different experiments the GPU performance variation will be eliminated; therefore, it provides a standardized, reproducible benchmark.
-During our experiment, we will use two sets of $\{c_{i}\}$, existing $\{c_{i}\}$ reported by (Leviathan et al., 2023) and an heuristic $\{c_{i}\}$ being the ratio of number of tunable parameters between $\mathcal{M}_{di}$ and $\mathcal{M}_{t}$.
+The benefit of the $SWI$ evaluation is that $count(S,\mathcal{M}_{t},T)$ and $count(S,\mathcal{M}_{di},T)$ will be the same across different systems, so by using the same $c_{i}$ across different experiments the GPU performance variation will be eliminated; therefore, it provides a standardized, reproducible measurement.
+During our experiment, we will use two sets of $\{c_{i}\}$, existing $\{c_{i}\}$ reported by Leviathan et al.
+(2023) and an heuristic $\{c_{i}\}$ being the ratio of number of tunable parameters between $\mathcal{M}_{di}$ and $\mathcal{M}_{t}$.
 Dataset Algorithm $\{\mathcal{M}_{di}\}$ SWI (MS) $k_{11}$ $k_{21}$ $k_{12}$ $l$
@@
 GSM8K C.S. Drafting small, mag $3.16$ 11 - - -
-GSM8K C.S. Drafting base, small, mag 3.75 7 1 3 1.5
+GSM8K C.S. Drafting base, small, mag 3.75 7 1 10 1.5
 MMLU Autoregressive - 1 - - - -
@@
 MMLU C.S. Drafting small, mag $4.36$ 17 - - -
-MMLU C.S. Drafting base, small, mag 4.83 5 1 14 2
+MMLU C.S. Drafting base, small, mag 4.83 5 1 19 2
 Dataset Algorithm $\{\mathcal{M}_{di}\}$ SWI (PW) $k_{11}$ $k_{21}$ $k_{12}$ $l$
@@
 GSM8K C.S. Drafting small, mag $2.90$ 7 - - -
-GSM8K C.S. Drafting base, small, mag 3.29 5 1 4 1.5
+GSM8K C.S. Drafting base, small, mag 3.29 5 1 9 1.5
 MMLU Autoregressive - 1 - - - -
@@
 MMLU C.S. Drafting small, mag $3.97$ 12 - - -
-MMLU C.S. Drafting base, small, mag 4.23 5 1 8 1.5
-Table 2: The experiment results from dataset GSM8K and MMLU.
+MMLU C.S. Drafting base, small, mag 4.23 5 1 13 1.5
+Table 2: The experiment results from datasets GSM8K and MMLU.
 BASE and SMALL refer to FLAN-T5-base and FLAN-T5-small.
-SWI(model size) is the standardized walltime improvement with the assumption the latency of each run of a model is its number of parameters(MS) and time cost data from previous work(PW)(Leviathan et al., 2023) respectively.
+SWI (MS) is the standardized walltime improvement with the assumption that the latency of each run of a model is its number of parameters (model size).
+SWI (PW) is the SWI with the assumption that the latency of each run of a model is the time cost data reported from previous work (Leviathan et al., 2023).
+$k_{11}$, $k_{21}$, $k_{12}$, $l$ are the hyperparameters.
+$k_{1i}$ represents the step limitation target model and the draft models, $k_{21}$ is the step limitations between the first and second draft model, and $l$ is leniency as shown in algorithm 1.
+For speculative decoding, the $k_{11}$ is simply the $k$.
 6 Experiment
 6.1 Setup
-We choose our target and draft models from FLAN-T5 (Chung et al., 2022) family for our experiment as there are large variations in model sizes in the FLAN-T5 family (from 77 million to 11 billion); also, the reasoning ability of the FLAN-T5 models providing more options for our experiments.
-To make our experiment similar to real-world usage, We do not perform fine-tuning and the generation is done in a zero-shot manner.
-We use FLAN-T5-xxl as our target model, and T5-base and T5-small as our reviewing draft models and Max-Gram algorithm to be the generating draft model.
-With limited computational budge, we are only able to perform sampling in temperate 00 (greedy decoding) due to the extra steps of hyperparameter tuning needed for the Cascade Speculative Decoding algorithm; the reason behind this is the benefit of temperate 00 that we can reuse the target model output which is deterministic which saves a great amount of computational resource on running the target models repeatedly.
-Cascade Speculative Drafting Inference To simplify the number of hyperparameters to tune, we use MaG to generate 10 tokens at once as it’s rare that more than 10 tokens will be accepted, and we do not use lenience when the reviewer is $\mathcal{M}_{t}$ to ensure the output distribution remains do not change, and we also eliminate lenience between $\mathcal{M}_{d2}$ and MaG due to the performance map between them are large.
-With these constraints, we are left with at most four parameters $k_{11}$, $k_{21}$, $k_{12}$, $l$.
-For the step of CS Drafting when the target model is the reviewer $k_{11}$ and $k_{12}$ are used.
-$k_{21}$, $l$ in the step when $\mathcal{M}_{d1}$ is the reviewer.
-However, there are still too many hyperparameters for us to search for the optimal result exhaustively; instead, we use halving grid search to search for a set of hyperparameters to compare to the baseline.
-Baseline We consider speculative decoding and vanilla autoregressive generation as our baseline.
-For speculative decoding, we consider T5-small and T5-base as the draft model as they perform better than T5-large in previous work(Leviathan et al., 2023) and our preliminary experiment.
-We try a large range of k-values from 2 to 30 to ensure that we will find the best result for each model as the draft model to serve as our baseline.
-Dataset We chose reasoning tasks for our experiment as they present more challenges of real-world NLP applications than summarization tasks.
-For both datasets, we experiment in a zero-shot chain-of-thought (Wei et al., 2023) setup.
-The datasets we experimented on are the following:
-GSM8k (Cobbe et al., 2021) A challenging mathematical reasoning dataset many large language models struggle with.
-MMLU (Hendrycks et al., 2021) A reasoning dataset in many different subjects.
-MMLU does not have training sets meaning applying task-specific methods such as knowledge distillation to improve speculative decoding can be challenging.
+We chose our target and draft models from the FLAN-T5 (Chung et al., 2022) family for our experiment, as there is a large variation in model sizes within the FLAN-T5 family (ranging from 77 million to 11 billion parameters).
+To align our experiment with current common usage, we do not perform fine-tuning, and the generation is conducted in a zero-shot manner.
+We use FLAN-T5-xxl as our target model, FLAN-T5-base and FLAN-T5-small as our reviewing draft models, and the Max-Gram algorithm as the generating draft model.
+To ensure our experiments are fully reproducible, we perform sampling at temperature 00, which is essentially greedy decoding.
+And we do not observe any significant difference between sampling with temperature $1$ and greedy decoding in previous speculative decoding experiments (Leviathan et al., 2023).
+Cascade Speculative Drafting Inference To reduce the number of hyperparameters to tune, we use MaG to generate 10 tokens at once, as it is rare for more than 10 tokens to be accepted.
+We do not use lenience when the reviewer is $\mathcal{M}_{t}$ to ensure the output distribution remains does not change.
+Additionally, we eliminate leniency between $\mathcal{M}_{d2}$ and MaG because the performance gap between them is large.
+With these constraints, we are left with at most four hyperparameters: $k_{11}$, $k_{21}$, $k_{12}$, and $l$.
+For the CS Drafting step where the target model is the reviewer, $k_{11}$ and $k_{12}$ are used.
+$k_{21}$ and $l$ are used in the step where $\mathcal{M}_{d1}$ is the reviewer.
+However, there are still too many hyperparameters for us to exhaustively search for the optimal result; instead, we employ a halving grid search to find a set of hyperparameters.
+Baseline We consider both speculative decoding and vanilla autoregressive generation as our baselines.
+For speculative decoding, we opt for FLAN-T5-small and FLAN-T5-base as the draft models, as they perform better than FLAN-T5-large, according to Leviathan et al.
+(2023) and our preliminary experiments.
+We test a wide range of $k$-values, from 2 to 30, to ensure that we find the best result for each model serving as our draft model to establish our baseline.
+Dataset We chose two commonly used datasets for our experiments. For both datasets, we conducted experiments in a zero-shot chain-of-thought setup (Kojima et al., 2022; Wei et al., 2023):
+  • GSM8K (Cobbe et al., 2021) is a dataset comprising 8,500 high-quality, linguistically diverse, grade-school math word problems.
+It focuses on multi-step reasoning with problems that are typically solvable using basic arithmetic in 2-8 steps.
+  • MMLU (Hendrycks et al., 2021), or Massive Multitask Language Understanding, is a benchmark for testing how well large language models grasp knowledge.
+It encompasses 57 diverse subjects, ranging from elementary science to advanced law.
 6.2 Experimental Results
-The experiments are shown in table 2.
-In two settings of SWI and Walltime, Cascade Speculative Decoding has exceeded the speculative decoding algorithm.
-For GSM8K, Cascade Speculative Decoding has a maximum improvement of $37\%$ over the fastest speculative algorithm; for MMLU, the maximum improvement of Cascade Speculative Decoding over speculative decoding is $72\%$.
-One other observation is that despite during speculative decoding t5-small has out-performed t5-base on some experiments, but in CS Drafting with only one neural model and MaG, T5-base consistently outperforms t5-small.
+Table 2 presents the experimental results.
+In two settings of SWI, Cascade Speculative Decoding has outperformed the speculative decoding algorithm.
+For GSM8K, Cascade Speculative Decoding achieved a maximum additional speedup of 37% over the fastest speculative algorithm; for MMLU, the maximum additional speedup improvement over speculative decoding is 72%.
+Additionally, we emphasize the effectiveness of the MaG algorithm.
+When comparing CS Drafting with one neural model and MaG against the fastest speculative decoding setup, we found that CS Drafting with one neural model gained up to a 56% speedup on MMLU and a 24% speedup on GSM8K.
+Notably, the MaG algorithm only involves a bigram model with as many parameters as the size of the tokenizer, making its memory cost negligible.
+Thus, the speedup gained using CS Drafting with one neural model involves no additional deployment overhead while reducing both latency and computational cost, making it a superior choice over speculative decoding.
+When comparing CS Drafting with two neural models to CS Drafting with one neural model, the former requires additional memory utilization.
+Despite the latency reduction, CS Drafting with two neural models can potentially incur more computational cost.
+However, during our experiments, since the draft models had only 2% of the number of parameters of the target model, the memory cost did not outweigh the reduced latency.
+Therefore, the CS Drafting algorithm with two neural models still uses less computational resources, as it completes each generation round faster.
+Another observation is that despite FLAN-T5-small sometimes outperforming FLAN-T5-base as a draft model for speculative decoding, in CS Drafting with the aid of MaG, FLAN-T5-base consistently outperforms FLAN-T5-small.
+This raises a topic for future study: determining the ideal size of draft models for CS Drafting.
 7 Related Work
-7.1 Efficienct Methods for Natural Language Inference
+7.1 Efficienct Methods for Language Model Inference
 In the era of large language models, the efficiency during inference becomes a key to model service.
@@
 8 Conclusion
-In our work, we proposed a novel algorithm CS.
-Drafting that involves two cascades, the vertical cascade, and the horizontal cascade.
-The vertical cascade eliminates the necessity of autoregressive generation from a neural language model, and the horizontal cascade effectively allocates the cost for drafting tokens at different positions.
-Our theories proved the optimality of the horizontal cascade, and empirically, we showed that CS.
-Drafting achieved up to 72 percent more speedup over speculative decoding without any additional training while maintaining generation quality.
+In our work, we propose a novel algorithm, CS Drafting, which involves two cascades: the vertical cascade and the horizontal cascade.
+The vertical cascade eliminates the necessity of autoregressive generation from a neural language model, while the horizontal cascade effectively allocates the cost of drafting tokens at different positions.
+Our experiments demonstrate that CS Drafting achieves up to a 72 percent additional speedup over speculative decoding without any additional training, while maintaining the same generation quality as the target model.
 Appendix A Max-Gram Implementation
@@
 3
-4def max_gram(input_ids, encoder_ids, n=1):
+4
+5def max_gram(input_ids, encoder_ids, n=1):
 6    matches = (encoder_ids[0] == input_ids[0, -1]).int()
@@
 $=\frac{f(r)(\alpha_{d_{1}}A_{1}+A_{2})+(1-f(r))(\alpha_{d_{2}}A_{1}+A_{2})}{f(r)(1+c_{d_{1}}+\sum_{i=1,i\neq r}^{k}c_{i})+(1-f(r))(1+c_{d_{2}}+\sum_{i=1,i\neq r}^{k}c_{i})}$
-$\leq\max{\frac{(\alpha_{d_{1}}A_{1}+A_{2})}{1+c_{d_{1}}+\sum_{i=1,i\neq r}^{k}c_{i}},\frac{\alpha_{d_{2}}A_{1}+A_{2})}{1+c_{d_{2}}+\sum_{i=1,i\neq r}^{k}c_{i})}}$
-Meaning that $f(r)=0$ or $f(r)=1$ will result in a time improvement factor no less than a non-integer $f$, meaning replacing $f(r)$ with an integer will still result in an optimal system.
-This contradicts with the assumption that $f$ is the system with the largest $r$.
-As a result, there exists an optimal system without interpolation.
+$\leq\max{(\frac{\alpha_{d_{1}}A_{1}+A_{2}}{1+c_{d_{1}}+\sum_{i=1,i\neq r}^{k}c_{i}},\frac{\alpha_{d_{2}}A_{1}+A_{2}}{1+c_{d_{2}}+\sum_{i=1,i\neq r}^{k}c_{i}})}$
+This means that $f(r)=0$ or $f(r)=1$ will result in a time improvement factor that is no less than that of a non-integer $f$, indicating that replacing $f(r)$ with an integer will still yield an optimal system.
+This contradicts the assumption that $f$ is the system with the largest $r$.
+As a result, there exists an optimal system that does not require interpolation.
 ∎

### v3

@@
 Abstract
-Speculative decoding reduces the latency of large language model (LLM) inference by leveraging a draft model to draft for a larger target model to review.
-However, drafting in speculative decoding involves slow autoregressive generation and allocates the same amount of time to generating tokens of varying importance.
-These two inefficiencies lead to its suboptimal performance.
-To address this issue, we introduce Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that comprises two types of cascades.
-The Vertical Cascade eliminates autoregressive generation from neural models, while the Horizontal Cascade constitutes efficient time allocation in drafting.
-Combining both cascades, our CS Drafting algorithm has achieved up to a 72 percent additional speedup over speculative decoding in our experiments, while maintaining the same output distribution as the target model.
+Introduced to enhance the efficiency of large language model (LLM) inference, speculative decoding operates by having a smaller model generate a draft.
+A larger target model then reviews this draft to align with its output, and any acceptance by the target model results in a reduction of the number of the target model runs, ultimately improving efficiency.
+However, the drafting process in speculative decoding includes slow autoregressive generation and allocates equal time to generating tokens, irrespective of their importance.
+These inefficiencies collectively contribute to the suboptimal performance of speculative decoding.
+To further improve LLM inference, we introduce Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that incorporates two types of cascades.
+The Vertical Cascade eliminates autoregressive generation from neural models, while the Horizontal Cascade optimizes time allocation in drafting for improved efficiency.
+Combining both cascades, CS Drafting achieves up to an 81 percent additional speedup over speculative decoding in our experiments, while maintaining the same output distribution as the target model.
+Our code is publicly available at https://github.com/lfsszd/CS-Drafting.
 Machine Learning, ICML
@@
 1 Introduction
-The advent of Large Language Models (LLMs), like GPT-4 (OpenAI, 2023) has marked a significant milestone in the field of natural language processing (NLP).
-These models have not only excelled in various NLP tasks but have also found widespread application in user-interactive settings such as chatbots and virtual assistants.
-This application involves an extremely high number of users, up to hundreds of millions daily.
+The advent of Large Language Models (LLMs), like GPT-4 (OpenAI, 2023), has marked a significant milestone in the field of natural language processing (NLP).
+These models have not only excelled in various NLP tasks but have also found widespread applications in user-interactive settings, such as chatbots and virtual assistants.
+However, these applications involve an extremely high number of users, up to hundreds of millions daily.
 To serve in real-time at this scale, a low-latency system is not only cost-saving but also crucial for keeping the service running.
-In addition, the sheer scale of the service means that even a slight improvement in the latency of LLMs can greatly contribute to both the provider and the community.
-Consequently, optimizing LLM latency has become a crucial and impactful research focus.
-Unfortunately, LLMs currently suffer from high latency, especially in long-form generation, as autoregressive LLMs generate tokens one by one.
-An emerging solution, known as speculative decoding (Leviathan et al., 2023; Chen et al., 2023), shows potential to mitigate this issue.
-In speculative decoding, a draft model (which is smaller and faster) generates $k$ tokens in each step (with $k$ being a hyperparameter), and these tokens are then reviewed by a target model (which is larger and slower) in parallel.
-The target model will accept any tokens aligned with its output and further generate one token; the process happens all in a single run while maintaining the output distribution identical to the autoregressive one by the target model itself.
-As a larger number of expected accepted tokens per run indicates fewer target model runs, the efficiency of speculative decoding can be attributed to two main factors: the expected number of tokens accepted by the target model per run, and the latency of drafting.
-Figure 1: The probability of acceptance of draft tokens in relation to their positions in a single step of speculative decoding, evaluated on FLAN-T5-small, base, and large models.
+In addition, the sheer scale of the service means that even a slight improvement in the latency of LLMs can greatly contribute to both the service provider and the community.
+Consequently, optimizing the latency of LLMs has become a critical area of research.
+Unfortunately, the ever-growing size of LLMs significantly increases the latency, especially in long-form generation, as autoregressive LLMs generate tokens one by one.
+An emerging solution, known as speculative decoding (Leviathan et al., 2023; Chen et al., 2023; Xia et al., 2023), shows potential to mitigate this issue.
+In speculative decoding, a draft model (which is smaller and faster) generates $k$ tokens in each step (with $k$ being a hyperparameter) autoregressively, and these tokens are then reviewed by a target model (which is larger and slower) in parallel.
+In one single run, the target model will accept any tokens aligned with its output and further generate one token.
+The drafting process in speculative decoding enables the target model to generate multiple tokens in a single run while maintaining its output distribution unchanged.
+With a properly sized draft model, speculative decoding achieves a speedup of 2 to 3 times, making it a potential method for solving high latency issues.
+Figure 1: The figure shows an example of Cascade Speculative Drafting with target model $\mathcal{M}_{t}$ and draft models $\mathcal{M}_{d_{1}}$, $\mathcal{M}_{d_{2}}$, and $\mathcal{M}_{d_{3}}$.
+The horizontal cascade involves using larger draft models to generate the earlier tokens and smaller models to generate the later tokens.
+The vertical cascade requires each model to review drafts from smaller models with the exception of the smallest model which is a statistical language model.
+As the horizontal cascade and vertical cascade are orthogonal to each other, CS Drafting combines both approaches for optimal efficiency.
+However, since draft models are typically required to generate multiple tokens in multiple steps, where each generation still involves inefficient autoregressive decoding, the performance of speculative decoding could be limited by the drafting latency.
+This inefficiency is also indicated by Leviathan et al.
+(2023)’s experiments, where it was observed that very small models (e.g., around two orders of magnitude smaller than the target model) are usually the best choice for drafting because their inference cost is lower compared to that of a larger draft model, despite the fact that larger draft models usually have higher-quality generation.
+This underscores that improving drafting efficiency is crucial for further enhancing the performance of speculative decoding.
+In light of this, one key strategy to address this bottleneck is to avoid the inefficient autoregressive generation of neural draft models.
+Based on this consideration, it is noted that statistical language models, such as bigram language models, incur negligible latency and computational resource costs compared to neural language models, owing to their simple structure.
+However, because the tokens generated by statistical language models usually do not have a high probability of being accepted by the target model, speculative decoding with statistical language models alone may not yield optimal results compared to using a well-sized neural language model from the same family as the draft model.
+Nonetheless, we notice that it is not necessary to use only one draft model in speculative decoding—statistical language models can serve as the “draft” model for the neural draft model, thereby eliminating autoregressive generation from the neural draft model.
+Figure 2: The probability of acceptance of draft tokens in relation to their positions in a single step of speculative decoding, evaluated on FLAN-T5-small, base, and large models on GSM8K.
 The draft model generates 30 tokens at each step.
-Figure 2: Cascade Speculative Sampling uses LLMs of multiple sizes to generate and verify tokens.
-The model at the inner layer $\mathcal{M}_{d(i)}$ generates tokens proposed to and verified by LLMs at the next outer layer, $\mathcal{M}_{d(i+1)}$.
-This recursive process forms the vertical cascade.
-After a vertical cascade, a smaller $\mathcal{M}_{d(i)}$ will serve as the target model and perform a vertical cascade.
-The process repeats until the smaller model becomes a statistical language model that will generate autoregressively forming the horizontal cascade.
-[…]
-In light of this, there exist a type of draft models, including statistical language models like bigram language models, that are immune to drafting efficiency issues.
-Owing to their simple structure, these models incur negligible latency and computational resource costs compared to neural language models, thereby avoiding the aforementioned trade-off.
-Therefore, assuming enough parallel computing power to perform computation on additional $k$ tokens with similar time cost (Leviathan et al., 2023), utilizing speculative decoding with a statistical language model as the draft model can generally and intuitively enhance the efficiency of a neural language model.
-However, due to their limited capabilities, speculative decoding with statistical language models alone may not yield optimal results compared to using a well-sized neural language model from the same family as the target model when the neural language model is available.
-Based on the above analysis, we propose Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that comprises multiple draft models, starting with the smallest being a statistical language model.
+The results on MMLU are shown in Figure 3 of the Appendix.
+Furthermore, our analysis in Figure 2 reveals a pattern during the drafting step: tokens generated later in the sequence by the draft model show a progressively lower probability of being accepted by the target model.
+This is because the probability of a token being accepted is conditioned on the acceptance of the previous tokens.
+It indicates that later tokens from draft models are more prone to rejection, contributing less to the expected number of accepted tokens per draft step, yet incurring the same latency.
+Inspired by the above observations, we propose Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that comprises multiple draft models, starting with the smallest being a statistical language model.
 Each neural draft model reviews generations from a smaller model and then proposes its reviewed content to either a larger draft model or the target model.
+In this design, the drafting of each neural model will be accelerated by drafting from a smaller model, avoiding the inefficiency of autoregressive generation from neural models.
 We refer to this tiered speculative decoding approach as the Vertical Cascade.
-Furthermore, our analysis in Figure 1 reveals an interesting pattern in the expected acceptance rate during the drafting step: tokens generated later in the sequence by the draft model show a progressively lower probability of being accepted by the target model [^1] A theoretical analysis is provided in Section 4.
-Note that this probability differs from the acceptance rate, which is the probability of a token being accepted conditioned on the previous tokens’ acceptance..
-This indicates that later tokens from draft models, more prone to rejection, contribute less to the expected number of accepted tokens, yet incur the same latency.
-Based on this observation and our practice, we suggest the use of smaller, faster draft models for generating these high-rejection tokens, forming the Horizontal Cascade.
+In addition, we suggest the use of smaller, faster draft models for generating high-rejection tokens that are trailing in drafting generation, forming the Horizontal Cascade.
 Along with the aforementioned Vertical Cascade, these strategies compose our complete CS Drafting approach, as illustrated in Figure 1.
-Through theoretical analysis and empirical studies, we demonstrate that the CS Drafting algorithm outperforms speculative decoding in terms of latency across various tasks and settings, achieving an additional speedup of up to 72% over speculative decoding.
+Through theoretical analysis and empirical studies, we demonstrate that the CS Drafting algorithm outperforms speculative decoding in terms of latency across various tasks and settings, achieving an additional speedup of up to 81% over speculative decoding.
 These findings highlight the practical advantages and efficiency enhancements offered by both vertical and horizontal cascades.
-Our main contributions are summarized as follows:
+The main contributions are summarized as follows:
   • We introduce Cascade Speculative Drafting (CS Drafting), a speculative-execution-based algorithm that improves language model inference speed without sacrificing generation quality.
@@
 2 Preliminary
-2.1 Speculative Decoding Algorithm
-In real-world applications of large language models (LLMs), the sequential nature of autoregressive decoding emerges as a primary latency bottleneck in terms of response time.
-Addressing this, speculative decoding (SD) (Leviathan et al., 2023) presents an innovative solution that utilizes a compact draft model for sequential token generation with validation by a larger target model resulting in reduced latency.
-The core concept of speculative decoding involves a novel algorithm that accelerates sampling from autoregressive models without altering output distributions.
+The core concept of speculative decoding (Leviathan et al., 2023) involves the utilization of a small draft model for sequential token generation with validation by a larger target model resulting in reduced latency.
+This design accelerates sampling from autoregressive models without altering output distributions.
 At its heart, there are two key observations: 1) certain generations in language modeling are simpler than others and can be predicted by more efficient models correctly, and 2) using speculative execution along with a new sampling method enables faster, exact decoding from large models.
-Specifically, let $x$ be the input tokens at a run and $\mathcal{M}_{t}$ and $\mathcal{M}_{d}$ are the target and the draft model respectively, and $\mathcal{M}_{t}(x)[i]$ and $\mathcal{M}_{d}(x)[i]$ be their probability output at $i$-th token when input is $x$.
+Specifically, let $x$ be the input tokens at a run and $\mathcal{M}_{t}$ and $\mathcal{M}_{d}$ are the target and the draft model respectively, $k$ be the number of draft tokens generated per step, and $\mathcal{M}_{t}(x)[i]$ and $\mathcal{M}_{d}(x)[i]$ be their probability output at $i$-th token when input is $x$.
 We interpret speculative sampling as a two-stage operation.
-In the proposing stage, we sample $\{x_{t+1},...,x_{t+k}\}$ from draft model $\mathcal{M}_{d}$ autoregressively and append them to $x$ In the reviewing stage, let $x_{i}\in\{x_{t+1},...,x_{t+k}\}$ represents the token at the current position, and we accept it if $\mathcal{M}_{d}(x)[i-1]<\mathcal{M}_{t}(x)[i-1]$; in the event that $\mathcal{M}_{d}(x)[i-1]>\mathcal{M}_{t}(x)[i-1]$, we reject $x_{i}$ with a probability of $1-\frac{\mathcal{M}_{t}(x)[i-1]}{\mathcal{M}_{d}(x)[i-1]}$ and proceed to resample $x_{i}$ from a recalibrated distribution $\text{norm}(\max(0,\mathcal{M}_{t}(x)[i-1]-\mathcal{M}_{d}(x)[i-1]))$ and reject any token following $x_{i}$.
+In the proposing stage, we sample $\{x_{t+1},...,x_{t+k}\}$ from draft model $\mathcal{M}_{d}$ autoregressively and append them to $x$.
+In the reviewing stage, let $x_{i}\in\{x_{t+1},...,x_{t+k}\}$ represents the token at the current position, and we accept it if $\mathcal{M}_{d}(x)[i-1]\leq\mathcal{M}_{t}(x)[i-1]$; in the event that $\mathcal{M}_{d}(x)[i-1]>\mathcal{M}_{t}(x)[i-1]$, we reject $x_{i}$ with a probability of $1-\frac{\mathcal{M}_{t}(x)[i-1]}{\mathcal{M}_{d}(x)[i-1]}$ and proceed to resample $x_{i}$ from a recalibrated distribution $norm(\max(0,\mathcal{M}_{t}(x)[i-1]-\mathcal{M}_{d}(x)[i-1]))$ and reject any token following $x_{i}$.
 At the end, the target model will generate one additional token following the accepted tokens.
-Such a design guarantees the output is the same as sampling autoregressively using the target model alone (Leviathan et al., 2023).
-This method was empirically validated on various tasks and model sizes, demonstrating a significant acceleration in inference times (2x-3x faster) compared to standard implementations, without affecting the outputs.
-Importantly, it does not require task-specific training, altering model architectures, or changing training procedures, making it a practical solution for enhancing the efficiency of LLMs.
-3 Cascaded Speculative Drafting
-Algorithm 1 CascadeSpeculativeDrafting
-  Require:draft model $\{\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{n}}\}$, target model $\mathcal{M}_{t}$, $prefix$, flag $isFirstCall$, hyperparameters $K_{nn}$, $l$
+Such a design guarantees the output is the same as sampling autoregressively using the target model alone, as proven in Leviathan et al.
+(2023).
+Speculative decoding was empirically validated on various tasks and model sizes, demonstrating a significant acceleration in inference times (2x-3x faster) compared to standard implementations, without affecting the outputs.
+Importantly, it does not require task-specific training, altering model architectures, or changing training procedures, making it a practical solution for reducing the latency of LLM inference.
+3 Cascade Speculative Drafting
+In this section, we introduce our proposed method, Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that incorporates two types of cascades: vertical cascade and horizontal cascade.
+Algorithm 1 CascadeSpeculativeDraftingStep
+  Require: draft models $\{\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{n}}\}$, target mode $\mathcal{M}_{t}$, $prefix$, flag $isFirstCall$, hyperparameters $K_{nn}$, $l$
   draftList $\leftarrow$ $[\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{n}}]$
-   $\triangleright$ Initialize curGen and curProb
-  curGen $\leftarrow$ $prefix$, curProbs $\leftarrow$ a list of ones of length prefix
-   $\triangleright$ Unpack the a list of $k$ for current function call
+   $\triangleright$ Initialize curGen and curProb.
+  curGen $\leftarrow$ $prefix$, curProbs $\leftarrow$ a list of ones with the same length as prefix
+   $\triangleright$ Unpack the a list of $k$ for the current function call.
   $[k_{1},...,k_{n-1}]\leftarrow$ first row of $K_{nn}$
-   $\triangleright$ Base case of the recursive call when the MaG generate
-  if draftList only contains one element then
+   $\triangleright$ Generate using MaG for the Base case of the recursive call.
+  if draftList is empty then
      $\mathcal{M}\leftarrow$ first element of draftList
      $res\leftarrow$ $\mathcal{M}(curGen)$
-      $\triangleright$ MaG generates greedily, so the logits will be 1 at the position of the generation token and 0 otherwise
      return $res.generation,res.logits$
   end if
-   $\triangleright$ Begin the cascade
+   $\triangleright$ Perform the horizontal cascade with the for loop.
   for $i\leftarrow 1$ to $n$ do
+      $\triangleright$ Prepare the arguments for the next recursive call.
      curTarget $\leftarrow$ the $i$-th item of draftList
      curDraftList $\leftarrow$ the sublist of draftList starting from index $i+1$
-     curK $\leftarrow$ the submatrix of $K_{nn}$ from with the top-left corner at $(i+1,j)$ extending to the bottom-right corner
-     while length of curGen - prefix.length is less than $k_{i}$ do
-        $[x_{1},..,x_{r}],[p_{1},p_{2},...,p_{r}]\leftarrow$ CascadeSpeculativeDecoding(curDraftList, curTarget, curGen, False, curK, $l$)
-        curGen $\leftarrow[x_{1},..,x_{r}]$
-         $\triangleright$ Do not overwrite the previous probabilities
-[…]
-        Add elements of $[p_{s},p_{2},...,p_{r}]$ to curProbs
+     curK $\leftarrow$ the submatrix of $K_{nn}$ from with the top-left corner at $(i+1,i+1)$ extending to the bottom-right corner curPrefix $\leftarrow$ curGen
+     while curGen.length - curPrefix.length is less than $k_{i}$ do
+        curPrefix $\leftarrow$ curGen
+         $\triangleright$ Perform the vertical cascade with the recursive call.
+        $[x_{1},..,x_{u}],[p_{1},p_{2},...,p_{v}]\leftarrow$ CascadeSpeculativeDraftingStep(curDraftList, curTarget, curPrefix, False, curK, $l$)
+        curGen $\leftarrow[x_{1},..,x_{u}]$
+        s $\leftarrow$ curProbs.length + 1
+        Add elements of $[p_{1},p_{2},...,p_{v}]$ to curProbs
      end while
   end for
-   $\triangleright$ No leniency when the target model reviews
+   $\triangleright$ Set lenience to $1$ when the original target model reviews.
   if isFirstCall then
@@
   end if
-   $\triangleright$ Review process by $\mathcal{M}_{t}$ to ensure the output distribution is the same as autoregressive generation
-  $[p_{1}^{\prime},p_{2}^{\prime},...,p_{m}^{\prime}]\leftarrow$ $\mathcal{M}_{t}(curGen)$
-  $index\leftarrow m$
-  for $i\leftarrow 1+prefix.length$ to $m-1$ do
-     Sample $r$ from a uniform distribution
-     if $r\leq min(1,\frac{lp_{i}^{\prime}}{p_{i}})$ then
-        $index\leftarrow i$ and exit the loop
-     end if
-  end for
-  Sample $x_{last}$ from $p_{index}^{\prime}-p_{index}$
-  return $[x_{1},...,x_{index-1}]+[x_{last}]$, $[p_{1}^{\prime},p_{2}^{\prime},...,p_{index}^{\prime}]$
+   $\triangleright$ Use $\mathcal{M}_{t}$ to review the draft generation.
+  $[x_{1},...,x_{out}],[p_{1}^{\prime},p_{2}^{\prime},...,p_{out}^{\prime}]$ = review($\mathcal{M}_{t}$, curGen, curProbs, l)
+  return $[x_{1},...,x_{out}],[p_{1}^{\prime},p_{2}^{\prime},...,p_{out}^{\prime}]$
 3.1 Vertical Cascade
-The speculative decoding algorithm relies on the autoregressive generation of a smaller draft model.
-However, despite its smaller size, the draft model must run $k$ times for each target model run, and the cost can still be significant.
-To further reduce latency in the drafting process, we propose using an even smaller model to assist in drafting and employing the original draft model to review the generation of this smaller model.
-This process can be done recursively until it reaches the smallest draft model.
-The smallest draft model can be a statistical language model, such as a bigram language model, where the cost is negligible compared to neural language models.
-We refer to this recursive process as Vertical Cascade.
-Furthermore, to further improve speed, we can introduce leniency, a hyperparameter in speculative decoding that loosens the review process by the target model, allowing for faster speed at the trade-off of potentially differing results from the target model (Leviathan et al., 2023).
-With leniency $l\in[1,\infty)$, the acceptance condition for token $x_{i}$ in speculative decoding becomes $\mathcal{M}{d}(x)[i]<l\times\mathcal{M}{t}(x)[i]$.
-Otherwise, with a probability of $1-\frac{l\times\mathcal{M}{t}(x)}{\mathcal{M}{d}(x)}$, we reject $x_{i}$ and any following tokens.
-While the reduced quality introduced by leniency in speculative decoding is generally undesirable in practice, it fortunately does not affect the final output in our vertical cascade approach—we can only apply leniency in the inner speculative decoding.
-As long as leniency is not applied to the target model, the final quality is always guaranteed by the target model.
+A notable inefficiency of the speculative decoding algorithm is the reliance on the autoregressive generation of a smaller draft model.
+Since the draft model must run $k$ times for each target model run, the cost can still be significant despite its smaller size.
+In light of this, we reduce the drafting inefficiency by using an even smaller model to assist in drafting and employing the original draft model to review the generation of this smaller model.
+In addition, since this process can be performed again on the draft model that drafts for the original model, we recursively perform this process until it reaches a statistical draft model that involves negligent cost, such as a bigram language model.
+In this approach, we expect each recursion step will reduce the drafting latency without altering the output distribution.
+We refer to this recursive speculative approach as Vertical Cascade.
+Additionally, we incorporate lenience, a hyperparameter that loosens the review process by the target model, allowing for faster speed at the trade-off of potentially differing results from the target model (Leviathan et al., 2023).
+Lenience can be adopted during sampling or greedy decoding with speculative decoding.
+Let lenience $l\in[1,\infty)$.
+When sampling, the acceptance condition for token $x_{i}$ is transformed to $\mathcal{M}_{d}(x)[i]\leq l\times\mathcal{M}_{t}(x)[i]$.
+If the acceptance condition is not satisfied, with a probability of $1-\frac{l\times\mathcal{M}_{t}(x)}{\mathcal{M}_{d}(x)}$, we reject $x_{i}$ and any following tokens.
+[^1] For simplicity, we assume the probability outputs are not standardized.
+We refer the readers to Leviathan et al.
+(2023) for the discussion on standardized sampling.
+When performing greedy decoding, the acceptance condition becomes deterministic and is simply either $argmax\mathcal{M}_{d}(x)[i]=argmax\mathcal{M}_{t}(x)[i]$ or $\mathcal{M}_{d}(x)[i]\leq l\times\mathcal{M}_{t}(x)[i]$.
+For the speculative decoding algorithm, the reduced quality introduced by lenience is generally undesirable.
+However, for the vertical cascade approach, lenience affects the final output only if it is applied when the target model reviews.
+Therefore, we can limit the application of lenience in the vertical cascade only when draft models review and do not apply lenience when the target model reviews.
+This can ensure the final output is not altered while further reducing latency.
 3.2 Horizontal Cascade
-As illustrated in Figure 1, another key observation is that during the drafting steps of speculative decoding, not all drafting tokens are created equal.
+Another key observation is that during the drafting steps of speculative decoding, not all drafting tokens are created equal, as illustrated in Figure 2.
 The first draft token is more likely to be accepted as it only depends on itself; the last token is rarely accepted, as it has a chance of being reviewed only if all preceding tokens are accepted.
-From another perspective, one could also say that the first draft token bears the responsibility for all succeeding tokens, while the last token bears none.
-In Section 4, we provide a theoretical analysis of this phenomenon.
-Based on this observation, we propose better allocation of the time cost of drafting by using larger models to generate the preceding draft tokens and smaller models for the succeeding tokens.
-This design allows more time to be allocated to more effective tokens, leading to a reduction in overall latency.
-We refer to this design as the Horizontal Cascade.
+From a theoretical perspective, assume the event of acceptance of each token being a Bernoulli distribution with probably $p$, the probability of $n$-th token being accepted is $p^{n}$, implying an exponential decrease of value for tokens generated later in the sequence.
+Inspired by this observation, we designed Horizontal Cascade, an approach that improves time allocation by draft token allocation.
+Horizontal Cascade assigns the largest draft model to perform the generation of the first draft token due to its highest output alignment with the target model, and it progressively uses a smaller as the new draft token to be generated is less likely to be accepted.
+This process stops after the smallest model, i.e., a statistical language model finishes.
+This design reduces the time cost of generating unimportant draft tokens with a costly draft model, leading to a reduction in overall latency.
 3.3 Max-Gram for Better Statistical Drafting
-A statistical language model is special to speculative decoding, as it incurs almost no cost when drafting for its target model.
+As both Vertical Cascade and Horizontal Cascade remark cascade toward faster draft models, a statistical language model, which is the basis of the cascade, becomes essential for the efficiency of both approaches.
 In our pursuit of a more effective statistical language model, we noticed a general pattern: in language model generation, some words and phrases from the input query frequently reappear in the generated content.
-Based on this observation, we designed the Max-Gram (MaG) algorithm.
+In light of this observation, we designed the Max-Gram (MaG) algorithm.
 It greedily identifies maximal matches between the initial input (or existing generation) and tokens from the end of the generation.
-In cases where there is no match, we resort to a bigram model based on the probability distribution of Wikipedia (chosen to maintain the generality of our experiment).
+In cases where there is no match, we resort to a bigram model based on the probability distribution of Wikipedia (chosen to maintain the generality).
 Surprisingly, despite its simplicity, this straightforward and general algorithm has proven highly effective in aligning with neural language model outputs.
@@
 3.4 Algorithm
-Combining the horizontal and vertical cascades, the algorithm is presented in Algorithm 1.
-At its center, there is a for-loop that performs the horizontal cascade, while the recursive call forms the vertical cascade.
-Notably, the last element of the draft models is the MaG model.
-The algorithm requires a two-dimensional hyperparameter, $K_{nn}$, with each row serving as the stop criteria for a layer of recursive calls.
-For simplicity, we assume the leniency parameter $l$ is universal for the algorithm, except when the target model is under review; thus, the algorithm can benefit from the speedup of leniency without altering the output distribution.
-4 Theoretical Analysis
-In this section, we present some theoretical analyses of the proposed Cascade Speculative Drafting algorithm.
-4.1 Expected Walltime Improvement
-To assist our theoretical analysts, we establish the notions.
+Combining the horizontal and vertical cascades, the algorithm of cascade speculative decoding is presented in Algorithm 1.
+At its center, the horizontal cascade is realized by the for loop, while the vertical cascade is implemented through recursive calls.
+Notably, the MaG model is incorporated as the smallest draft model to avoid autoregressive generation from a neural model.
+An example of CS Drafting is shown in Figure 1.
+The algorithm requires an upper-triangular hyperparameter, $K_{nn}$, with each row serving as the stop criteria for a layer of recursive calls.
+For simplicity, we assume the lenience parameter $l$ is universal for the algorithm, except when the target model is under review; thus, the algorithm can benefit from the speedup of lenience without altering the output distribution.
+4 Analysis
+We begin with some notions. Let $\mathcal{M}_{t}$ be the target model, $\mathcal{M}_{d}$ be the draft model, and $k$ be the number of draft tokens generated per step.
+Expected acceptance rate $\alpha(\mathcal{M}_{t},\mathcal{M}_{d})$ is the probability of draft generation by $\mathcal{M}_{d}$ being accepted by target model $\mathcal{M}_{t}$.
+Cost coefficient $c(\mathcal{M}_{t},\mathcal{M}_{d})$ is the ratio of time for a single run of draft model $\mathcal{M}_{d}$ over target model $\mathcal{M}_{t}$.
+Expected walltime improvement factor (EWIF) is the expected time improvement achieved by an algorithm under the i.i.d. assumption of token acceptance.
+Despite the simple setting of EWIF, Leviathan et al. (2023) have demonstrated that it aligns with the experimental results in most instances. Therefore, our analysis will concentrate on EWIF.
+4.1 Vertical Cascade
+We analyze EWIF of vertical cascade using generating functions, a well-studied topic in combinatorial mathematics (West, 2021).
+The properties of generating functions are useful in the recursion and evaluation process making the our final expression simple.
+We begin with the derivation of the probability generating function for speculative decoding.
+Theorem 4.1.
+For speculative decoding between $\mathcal{M}_{t}$ and $\mathcal{M}_{d}$, let ${p_{i}}$ be the probability of generating $i$ tokens.
+The probability generating function of ${p_{i}}$ satisfies the following equation:
+$\phi_{(\alpha,k)}(x)=1+(x-1)\frac{1-\alpha^{k+1}x^{k+1}}{(1-\alpha x)},$ (1)
+where $\alpha=\alpha(\mathcal{M}_{t},\mathcal{M}_{d})$.
+Proof in Appendix C.1.
+Corollary 4.2.
+The EWIF of speculative decoding is $\frac{\phi_{(\alpha,k)}^{\prime}(1)}{(ck+1)}=\frac{1-\alpha^{k+1}}{(1-\alpha)(ck+1)}$.
+We use the generating function to derive the EWIF of a vertical cascade and analyze the case involving two draft models, $\mathcal{M}_{d_{1}}$ and $\mathcal{M}_{d_{2}}$.
+Theorem 4.3.
+Assume $k$ to be the speculative decoding parameter between $\mathcal{M}_{d_{1}}$ and $\mathcal{M}_{d_{2}}$, and $n$ to be the number of steps $\mathcal{M}_{t}$ reviews. The EWIF by this system is
+$\frac{1-\alpha\phi^{n}(\alpha)}{(1-\alpha)(1+nc_{d_{1}}+nkc_{d_{2}})},$ (2)
+where $\phi(x)=\phi_{(\alpha(\mathcal{M}_{d_{1}},\mathcal{M}_{d_{2}}),k)}(x)$, $\alpha=\alpha(\mathcal{M}_{t},\mathcal{M}_{d})$, and $c_{d_{1}},c_{d_{2}}$ be $c(\mathcal{M}_{t},\mathcal{M}_{d_{1}}),c(\mathcal{M}_{t},\mathcal{M}_{d_{2}})$ respectively.
+Corollary 4.4.
+$\phi_{\alpha^{\prime},k}(\alpha)<\alpha$ for any $01>\alpha>0,1>\alpha^{\prime}>0,k>$ 0, so if $c_{d_{2}}<<1$, the EWIF of $\mathcal{M}_{d_{1}}$ and $\mathcal{M}_{d_{2}}$ is higher than EWIF of $\mathcal{M}_{d_{1}}$ alone.
+Proof in Appendix C.2.
+Therefore, with the statistical model having negligible cost (i.e., $c_{d_{2}}<<1$), it can almost always improve the efficiency of an SD system.
+4.2 Horizontal Cascade
+We also present an analysis of the walltime improvement offered by the horizontal cascade.
+To assist the analysis, we establish the notions.
 Let $\mathcal{M}_{t}$ be the target model, $\{\mathcal{M}_{i}\}$ be the draft models assisting generation with $\mathcal{M}_{i}$ being the draft model generating the $i$-th token.
@@
 Let $x$ be the input to the model at a single run, $\mathcal{M}_{t}(x)$ and $\mathcal{M}_{i}(x)$ are then the output probability distribution with input $x$.
-Expected acceptance rate $\alpha=\mathbf{E}(1-\frac{1}{2}|\mathcal{M}_{t}(x)-\mathcal{M}_{i}(x)|)$ is the probability of draft generation by $\{\mathcal{M}_{i}\}$ being accepted by target model $\mathcal{M}_{t}$ (Leviathan et al., 2023).
-Cost coefficient $c_{i}$ is the ratio of time for a single run of draft model $\mathcal{M}_{i}$ over target model $\mathcal{M}_{t}$.
-Expected walltime improvement factor (EWIF) is the expected time improvement of an algorithm based on. We provide the formula for EWIF in an extended setting when multiple draft models are involved.
+To simplify notation, let $\alpha_{i}=\alpha(\mathcal{M}_{t}(x),\mathcal{M}_{i}(x))$ and $c_{i}=c(\mathcal{M}_{t}(x),\mathcal{M}_{i}(x))$.
 Theorem 4.5.
-The expected walltime improvement factor (EWIF) by Algorithm 1 is $T(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})=\frac{\sum_{i=0}^{k}\prod_{j=1}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$.
-Proof.
-Since every token will only be considered by the target model if all tokens before are accepted, a simple recursion leads to the probability of acceptance of the $i$-th token being $\prod_{j=1}^{i-1}\alpha_{j}$, and the $\prod_{j=1}^{0}\alpha_{j}=1$ representing the token generated by the target model.
-In the right-hand side, the numerator is the expected number of tokens generated, and the denominator is the time costs for each step of such a system, so the right-hand side is the expected efficiency improvement.
-∎
-We can immediately achieve a formula for speculative decoding by requiring all the draft models to be the same.
-Corollary 4.2.
-When $q_{1}=q_{2}=...=q_{k}$, let $\alpha_{i}=\alpha$ and $c_{i}=c$ for any $i$, the EWIF is $\frac{1-\alpha^{k+1}}{(1-\alpha)(c\cdot k+1)}$.
+The expected walltime improvement factor (EWIF) of the horizontal cascade is $T(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})=\frac{\sum_{i=0}^{k}\prod_{j=1}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$.
 Furthermore, theorem 4.5 can be used to analyze the importance of the tokens in the drafting step.
@@
 Using the information provided by Leviathan et al.
-(2023), we calculat a simulated EWIF under the assumption that the event of acceptance by the target model is a Bernoulli trial.
+(2023), we calculate a simulated EWIF under the assumption that the event of acceptance by the target model is a Bernoulli trial.
 The results, shown in Table 1, indicate that speculative sampling with a horizontal cascade achieved better EWIF than vanilla speculative sampling under this assumption.
-4.2 Analysis of Model Interpolation
-The setting of cascade drafting can be simply extended using model interpolation that involves choosing between two models using randomness.
-Given two draft systems $d_{1}$ and $d_{2}$ (They can be two models generating autoregressive or two speculative decoding systems as used in vertical cascade) and a probability function $f$, an interpolation $d_{inter}$ is defined as $f(i)d_{1}+(1-f(i))d_{2}$.
-In other words, with probability $f(i)$, $d_{inter}$ is $d_{1}$, and with probability $1-f(i)$, $d_{inter}$ is $d_{2}$.
-As a result, $\alpha_{i}=f(i)\alpha_{d1}+(1-f(i))\alpha_{d2}=\alpha_{d2}+f(i)(\alpha_{d1}-\alpha_{d2})$ and $c_{i}=f(i)c_{d1}+(1-f(i))c_{d2}=c_{2}+f(i)(c_{d1}-c_{d2})$.
-An interpolation system can provide a smooth transition between two systems; however, in theorem 4.4, we prove that the optimal can always be reached without interpolation.
-Theorem 4.4.
-Given a horizontal cascade system with the setting mentioned above.
-There exists a probability function $f$ that optimizes the system such that $f$ satisfies $f(i)\in\{0,1\}$ for any $i$.
-This means each of the tokens will be generated by either $d_{1}$ or $d_{2}$, and the optimal system is deterministic.
-Proof of Theorem 4.4 is included in appendix B.
-Dataset $\mathcal{M}_{d1}$ $\mathcal{M}_{d2}$ $k_{1}$ $k_{2}$ EWIF
+Dataset $\mathcal{M}_{d_{1}}$ $\mathcal{M}_{d_{2}}$ $k_{1}$ $k_{2}$ EWIF
 CNNDM small - 9 - 2.65
@@
 In the simulation, speculative sampling with horizontal cascade exceeded the performance of the vanilla speculative decoding on both CNN Dailymail (Nallapati et al., 2016) and WMT EnDe (Bojar et al., 2018) datasets.
-5 Evaluation
+5 Experiments
+5.1 Evaluation Metric
 Previous works on speculative decoding and related methods relied on walltime as an evaluation method.
@@
 (2023) mentioned that cost coefficiency, the ratio between the time for a single run of draft model and target model will vary from system to system which can result in significant variation in wall time improvement across different systems, making comparisons across different research work difficult.
-In addition, a recent analysis suggests GPU speed can vary for the same GPU model with one being 1.5x faster than another GPU of the same model (Sinha et al., 2022), and even for the same GPU, the performance variation between different runs is unknown; therefore, theoretical walltime improvement up to 1.5x can be undermined by GPU variation, and methods without any improvement can be measured with an improvement up to 1.5x by running on a different GPU.
-The properties of such lack of standardization and stability make walltime unfit for measuring algorithmic improvement.
+In addition, a recent analysis suggests GPU speed can vary for the same GPU model with one being 1.5x faster than another GPU of the same model (Sinha et al., 2022); therefore, theoretical walltime improvement up to 1.5x can be undermined by GPU variation, and methods without any improvement can be measured with an improvement by running on a different GPU.
+The lack of standardization and stability makes walltime unfit for measuring algorithmic improvement.
 To resolve this issue, we design a new metric to measure speculative decoding and its related algorithms which can reduce the impact of varying/unstable GPU performance.
 Our proposed method, standardized walltime improvement (SWI), calculates the GPU times of the models, assuming that each run of a language model costs the same amount of time, an assumption made when inventing the speculative sampling algorithm (Chen et al., 2023).
-Given a speculative decoding system $S$ containing $k$ draft models $\mathcal{M}_{d1},...,\mathcal{M}_{dk}$, target model $\mathcal{M}_{t}$ solving a task $T$, and let $count(S,m,T)$ be the count of runs of model $m$ during the execution of task $T$ with system $S$, let $c_{i}$ be the expected cost coefficient being the time ratio between a single run of $\mathcal{M}_{i}$ and $\mathcal{M}_{t}$, and let $n$ be the total of tokens generated, the standardized walltime improvement is defined as
-$\begin{split}SWI&(\mathcal{M}_{d1},...,\mathcal{M}_{dk},\mathcal{M}_{t},T)=\\ &\frac{n}{count(S,\mathcal{M}_{t},T)+\sum_{i=1}^{n}count(S,\mathcal{M}_{di},T)c_{i}},\end{split}$
+Given a speculative decoding system $S$ containing $k$ draft models $\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{k}}$, target model $\mathcal{M}_{t}$ solving a task $T$, and let $count(S,m,T)$ be the count of runs of model $m$ during the execution of task $T$ with system $S$, let $c_{i}$ be the expected cost coefficient being the time ratio between a single run of $\mathcal{M}_{i}$ and $\mathcal{M}_{t}$, and let $n$ be the total of tokens generated, the standardized walltime improvement is defined as
+$\begin{split}SWI&(\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{k}},\mathcal{M}_{t},T)=\\ &\frac{n}{count(S,\mathcal{M}_{t},T)+\sum_{i=1}^{n}count(S,\mathcal{M}_{d_{i}},T)c_{i}},\end{split}$
 which is simply the sum of GPU time by all models when $c_{i}$ is measured during a local experiment.
-The benefit of the $SWI$ evaluation is that $count(S,\mathcal{M}_{t},T)$ and $count(S,\mathcal{M}_{di},T)$ will be the same across different systems, so by using the same $c_{i}$ across different experiments the GPU performance variation will be eliminated; therefore, it provides a standardized, reproducible measurement.
+The benefit of the $SWI$ evaluation is that $count(S,\mathcal{M}_{t},T)$ and $count(S,\mathcal{M}_{d_{i}},T)$ will be the consistent across different systems, so by using the same $c_{i}$ across different experiments the GPU performance variation will be eliminated; therefore, it provides a standardized, reproducible measurement.
 During our experiment, we will use two sets of $\{c_{i}\}$, existing $\{c_{i}\}$ reported by Leviathan et al.
-(2023) and an heuristic $\{c_{i}\}$ being the ratio of number of tunable parameters between $\mathcal{M}_{di}$ and $\mathcal{M}_{t}$.
-Dataset Algorithm $\{\mathcal{M}_{di}\}$ SWI (MS) $k_{11}$ $k_{21}$ $k_{12}$ $l$
+(2023) and an heuristic $\{c_{i}\}$ being the ratio of number of tunable parameters between $\mathcal{M}_{d_{i}}$ and $\mathcal{M}_{t}$.
+Dataset Algorithm $\{\mathcal{M}_{d_{i}}\}$ Speedup (MS) $k_{11}$ $k_{12}$ $k_{22}$ $l$
 GSM8K Autoregressive - 1 - - - -
-GSM8K S. Decoding base $3.38$ 10 - - -
-GSM8K S. Decoding small $3.06$ 11 - - -
-GSM8K C.S. Drafting base, mag $3.62$ 9 - - -
-GSM8K C.S. Drafting small, mag $3.16$ 11 - - -
-GSM8K C.S. Drafting base, small, mag 3.75 7 1 10 1.5
+GSM8K S Decoding base $3.38$ 10 - - -
+GSM8K S Decoding small $3.06$ 11 - - -
+GSM8K CS Drafting base, mag $3.70$ 10 - - -
+GSM8K CS Drafting small, mag $3.19$ 11 - - -
+GSM8K CS Drafting base, small, mag 3.88 8 13 1 3
 MMLU Autoregressive - 1 - - - -
-MMLU S. Decoding base $3.97$ 13 - - -
-MMLU S. Decoding small $4.12$ 19 - - -
-MMLU C.S. Drafting base, mag $4.50$ 12 - - -
-MMLU C.S. Drafting small, mag $4.36$ 17 - - -
-MMLU C.S. Drafting base, small, mag 4.83 5 1 19 2
-Dataset Algorithm $\{\mathcal{M}_{di}\}$ SWI (PW) $k_{11}$ $k_{21}$ $k_{12}$ $l$
+MMLU S Decoding base $3.97$ 13 - - -
+MMLU S Decoding small $4.12$ 19 - - -
+MMLU CS Drafting base, mag $4.56$ 13 - - -
+MMLU CS Drafting small, mag $4.39$ 14 - - -
+MMLU CS Drafting base, small, mag 4.88 5 19 1 5
+Dataset Algorithm $\{\mathcal{M}_{d_{i}}\}$ Speedup (PW) $k_{11}$ $k_{12}$ $k_{22}$ $l$
 GSM8K Autoregressive - 1 - - - -
-GSM8K S. Decoding base $2.99$ 8 - - -
-GSM8K S. Decoding small $2.76$ 8 - - -
-GSM8K C.S. Drafting base, mag $3.22$ 8 - - -
-GSM8K C.S. Drafting small, mag $2.90$ 7 - - -
-GSM8K C.S. Drafting base, small, mag 3.29 5 1 9 1.5
+GSM8K S Decoding base $2.99$ 8 - - -
+GSM8K S Decoding small $2.76$ 8 - - -
+GSM8K CS Drafting base, mag $3.27$ 9 - - -
+GSM8K CS Drafting small, mag $2.82$ 11 - - -
+GSM8K CS Drafting base, small, mag 3.43 5 9 1 3
 MMLU Autoregressive - 1 - - - -
-MMLU S. Decoding base $3.42$ 10 - - -
-MMLU S. Decoding small $3.51$ 11 - - -
-MMLU C.S. Drafting base, mag $4.07$ 9 - - -
-MMLU C.S. Drafting small, mag $3.97$ 12 - - -
-MMLU C.S. Drafting base, small, mag 4.23 5 1 13 1.5
-Table 2: The experiment results from datasets GSM8K and MMLU.
-BASE and SMALL refer to FLAN-T5-base and FLAN-T5-small.
-SWI (MS) is the standardized walltime improvement with the assumption that the latency of each run of a model is its number of parameters (model size).
-SWI (PW) is the SWI with the assumption that the latency of each run of a model is the time cost data reported from previous work (Leviathan et al., 2023).
-$k_{11}$, $k_{21}$, $k_{12}$, $l$ are the hyperparameters.
-$k_{1i}$ represents the step limitation target model and the draft models, $k_{21}$ is the step limitations between the first and second draft model, and $l$ is leniency as shown in algorithm 1.
+MMLU S Decoding base $3.42$ 10 - - -
+MMLU S Decoding small $3.51$ 11 - - -
+MMLU CS Drafting base, mag $4.21$ 6 - - -
+MMLU CS Drafting small, mag $3.99$ 13 - - -
+MMLU CS Drafting base, small, mag 4.32 5 8 1 5
+Table 2: The experimental results on FLAN-T5.
+Speedup (MS) is the standardized walltime improvement with the assumption that the latency of each run of a model is its number of parameters (model size).
+Speedup (PW) is the SWI with the assumption that the latency of each run of a model is the time cost data reported from previous work (Leviathan et al., 2023).
+$k_{11}$, $k_{12}$, $k_{22}$, $l$ are the hyperparameters.
+$k_{11}$ and $k_{12}$ represent the step limitation target model and the draft models, $k_{22}$ is the step limitations between the first and second draft model, and $l$ is lenience as shown in algorithm 1.
 For speculative decoding, the $k_{11}$ is simply the $k$.
-6 Experiment
 5.2 Setup
-We chose our target and draft models from the FLAN-T5 (Chung et al., 2022) family for our experiment, as there is a large variation in model sizes within the FLAN-T5 family (ranging from 77 million to 11 billion parameters).
+To ensure the generality of our findings, we perform experiments on both encoder-decoder and decoder-only models.
+For encoder-decoder models, we choose our target and draft models from the FLAN-T5 (Chung et al., 2022) family for our experiment, as there is a large variation in model sizes within the FLAN-T5 family (ranging from 77 million to 11 billion parameters).
+We use FLAN-T5-xxl as our target model, FLAN-T5-base and FLAN-T5-small as our reviewing draft models.
+For decoder-only models, we select LLAMA-2-chat-7B (Touvron et al., 2023) as the target model.
+Since there is no official release of a smaller model in LLAMA family, we use a pretrained 160M model with the same tokenizer as the reviewing draft model (Miao et al., 2024).
+In both cases, the Max-Gram algorithm is used as the generating draft model.
 To align our experiment with current common usage, we do not perform fine-tuning, and the generation is conducted in a zero-shot manner.
-We use FLAN-T5-xxl as our target model, FLAN-T5-base and FLAN-T5-small as our reviewing draft models, and the Max-Gram algorithm as the generating draft model.
-To ensure our experiments are fully reproducible, we perform sampling at temperature 00, which is essentially greedy decoding.
-And we do not observe any significant difference between sampling with temperature $1$ and greedy decoding in previous speculative decoding experiments (Leviathan et al., 2023).
+Since we do not observe any significant difference between sampling with temperature $1$ and greedy decoding in previous speculative decoding experiments (Leviathan et al., 2023), and to ensure our experiments are fully reproducible, we perform sampling at temperature 00, i.e., using greedy decoding by default.
 Cascade Speculative Drafting Inference To reduce the number of hyperparameters to tune, we use MaG to generate 10 tokens at once, as it is rare for more than 10 tokens to be accepted.
-We do not use lenience when the reviewer is $\mathcal{M}_{t}$ to ensure the output distribution remains does not change.
-Additionally, we eliminate leniency between $\mathcal{M}_{d2}$ and MaG because the performance gap between them is large.
-With these constraints, we are left with at most four hyperparameters: $k_{11}$, $k_{21}$, $k_{12}$, and $l$.
+We do not use lenience when the reviewer is $\mathcal{M}_{t}$ to ensure the output distribution does not change.
+We also avoid lenience between MaG and its reviewer, since there is still a significant performance gap between MaG and a neural model.
+With these constraints, we are left with at most four hyperparameters: $k_{11}$, $k_{12}$, $k_{22}$, and $l$.
 For the CS Drafting step where the target model is the reviewer, $k_{11}$ and $k_{12}$ are used.
-$k_{21}$ and $l$ are used in the step where $\mathcal{M}_{d1}$ is the reviewer.
-However, there are still too many hyperparameters for us to exhaustively search for the optimal result; instead, we employ a halving grid search to find a set of hyperparameters.
-Baseline We consider both speculative decoding and vanilla autoregressive generation as our baselines.
-For speculative decoding, we opt for FLAN-T5-small and FLAN-T5-base as the draft models, as they perform better than FLAN-T5-large, according to Leviathan et al.
+$k_{21}$ and $l$ are used in the step where $\mathcal{M}_{d_{1}}$ is the reviewer.
+Baselines We consider both speculative decoding and vanilla autoregressive generation as our baselines.
+When performing speculative decoding with the target model as FLAN-T5-xxl, we opt for FLAN-T5-small and FLAN-T5-base as the draft models, as they perform better than FLAN-T5-large, according to Leviathan et al.
 (2023) and our preliminary experiments.
+When LLAMA-2-chat-7B is the target model, we use the 160M model as the draft model.
 We test a wide range of $k$-values, from 2 to 30, to ensure that we find the best result for each model serving as our draft model to establish our baseline.
@@
 5.3 Experimental Results
-Table 2 presents the experimental results.
+Comparison to Baselines Table 2 presents the experimental results.
 In two settings of SWI, Cascade Speculative Decoding has outperformed the speculative decoding algorithm.
-For GSM8K, Cascade Speculative Decoding achieved a maximum additional speedup of 37% over the fastest speculative algorithm; for MMLU, the maximum additional speedup improvement over speculative decoding is 72%.
-Additionally, we emphasize the effectiveness of the MaG algorithm.
-When comparing CS Drafting with one neural model and MaG against the fastest speculative decoding setup, we found that CS Drafting with one neural model gained up to a 56% speedup on MMLU and a 24% speedup on GSM8K.
+For GSM8K, Cascade Speculative Decoding achieved a maximum additional speedup of 44% over the fastest speculative algorithm; for MMLU, the maximum additional speedup improvement over speculative decoding is 81%.
+Overall, CS Decoding achieves a 3-4x speedup over autoregressive generation, establishing itself as a faster algorithm over the baselines.
+Effectiveness of MaG When comparing CS Drafting with one neural model and MaG against the fastest speculative decoding setup, we found that CS Drafting with one neural model gained up to a 70% speedup on MMLU and a 32% speedup on GSM8K.
 Notably, the MaG algorithm only involves a bigram model with as many parameters as the size of the tokenizer, making its memory cost negligible.
-Thus, the speedup gained using CS Drafting with one neural model involves no additional deployment overhead while reducing both latency and computational cost, making it a superior choice over speculative decoding.
-When comparing CS Drafting with two neural models to CS Drafting with one neural model, the former requires additional memory utilization.
-Despite the latency reduction, CS Drafting with two neural models can potentially incur more computational cost.
-However, during our experiments, since the draft models had only 2% of the number of parameters of the target model, the memory cost did not outweigh the reduced latency.
-Therefore, the CS Drafting algorithm with two neural models still uses less computational resources, as it completes each generation round faster.
-Another observation is that despite FLAN-T5-small sometimes outperforming FLAN-T5-base as a draft model for speculative decoding, in CS Drafting with the aid of MaG, FLAN-T5-base consistently outperforms FLAN-T5-small.
-This raises a topic for future study: determining the ideal size of draft models for CS Drafting.
+Notably, the speedup gained using CS Drafting with one neural model involves no additional deployment overhead while reducing both latency and computational cost, making it a superior choice over speculative decoding.
+Draft Model Size Despite FLAN-T5-small mostly outperforming FLAN-T5-base as a draft model for speculative decoding, in CS Drafting with the aid of MaG, FLAN-T5-base consistently outperforms FLAN-T5-small.
+This implies that with the limitation of a single draft model, the ideal size of the draft model might increase with the assistance of the MaG model.
+Dataset Algorithm $\{\mathcal{M}_{d_{i}}\}$ Speedup $k_{11}$
+GSM8K Auto. - 1 -
+GSM8K S Decoding 160m $2.48$ 12
+GSM8K CS Drafting 160m, mag 2.86 15
+MMLU Auto. - 1 -
+MMLU S Decoding 160m $2.12$ 10
+MMLU CS Drafting 160m, mag 2.64 13
+Table 3: The experimental results on LLAMA.
+Results of Decoder-only Models As shown in Table 3, the results for decoder-only models exhibit similar traits to those observed in the experiments on encoder-decoder models.
+Therefore, for speculative decoding-based algorithms, we believe there is no distinct difference between an encoder-decoder model and a decoder-only model, aligning with the results in Leviathan et al.
+(2023).
 6 Related Work
 6.1 Efficienct Methods for Language Model Inference
-In the era of large language models, the efficiency during inference becomes a key to model service.
+In the era of large language models, efficiency during inference becomes a key to model service.
 To reduce the model inference cost and speed up, several efficient methods have been proposed, including pruning, knowledge distillation and quantization (Treviso et al., 2023).
@@
 6.2 Speculative Decoding
-Recent works improved the speed of speculative decoding by experimenting with task-specific knowledge distillation (Zhou et al., 2023), or using approaches similar to speculative decoding to achieve a tradeoff between quality and speed (Leviathan et al., 2023; Kim et al., 2023).
-However, the reduced quality for speed can be undesirable, and task-specific training might not be applicable to many real-world NLP applications such as chatbots; overall, we found no generic improvement for speculative decoding in the current literature.
 Speculative decoding (Chen et al., 2023; Leviathan et al., 2023) is proposed to improve inference speed by using a draft model to perform auto-regressive decoding and a target model to check and replace the generated tokens.
@@
 On the other hand, He et al.
-(2023) proposed constructing a datastore from a domain-specific corpus to serve as a draft model which can be used when a draft model is not available.
+(2023) propose constructing a datastore from a domain-specific corpus to serve as a draft model which can be used when a draft model is not available.
 7 Conclusion
-In our work, we propose a novel algorithm, CS Drafting, which involves two cascades: the vertical cascade and the horizontal cascade.
+In this work, we propose a novel algorithm, CS Drafting, which involves two cascades: the vertical cascade and the horizontal cascade.
 The vertical cascade eliminates the necessity of autoregressive generation from a neural language model, while the horizontal cascade effectively allocates the cost of drafting tokens at different positions.
-Our experiments demonstrate that CS Drafting achieves up to a 72 percent additional speedup over speculative decoding without any additional training, while maintaining the same generation quality as the target model.
+Our experiments show that CS Drafting achieves up to an 81 percent additional speedup over speculative decoding, while maintaining the same output distribution as the target model.
+Our work demonstrates that LLM inference can be further sped up by cascades without sacrificing generation quality or requiring additional task-specific training.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
 Appendix A Max-Gram Implementation
@@
 26    return encoder_ids[:, index+1:index + n+1]
-Appendix B Proof and Analysis of Theorem 4.4
-We first prove a lemma that helps the proof of theorem 4.4
-Lemma B.1.
-$\forall a,b,c,d\in\mathbb{R}_{>0},\gamma\in(0,1)$
-$\frac{\gamma a+(1-\gamma)b}{\gamma c+(1-\gamma)d}\leq\max(\frac{a}{c},\frac{b}{d})$
+Appendix B Additonal Results
+Figure 3: The probability of acceptance of draft tokens in relation to their positions in a single step of speculative decoding, evaluated on FLAN-T5-small, base, and large models on MMLU.
+Appendix C Proof
+C.1 Proof for Theorem 4.1
 Proof.
-Without loss of generality, assume $\frac{a}{c}\geq\frac{b}{d}$,
-$\frac{a}{c}\geq\frac{\gamma a+(1-\gamma)b}{\gamma c+(1-\gamma)d}$
-$\Longleftrightarrow a(\gamma c+(1-\gamma)d)\geq c(\gamma a+(1-\gamma)b)$
-$\Longleftrightarrow(1-\gamma)ad\geq(1-\gamma)cb$
-$\Longleftrightarrow\frac{a}{c}\geq\frac{b}{d}$
-Proof of Theorem 4.4
+The probability of accepting $i$ tokens is $\alpha^{i}-\alpha^{i+1}$, with the exception of the $k+1$-th token, which has a probability of $\alpha^{k}$ of being accepted.
+This is because it requires all the first $i$ tokens to be accepted and the $i+1$-th token to be rejected for this to happen.
+Therefore,
+$\phi_{(\alpha,k)}(x)=\alpha^{k}x^{k+1}+\sum_{i=0}^{k-1}(\alpha^{i}-\alpha^{i+1})x^{i+1}.$ (3)
+By rearranging the terms, we can achieve an expression much easier to work with
+$\phi_{(\alpha,k)}(x)=x+\sum_{i=1}^{k}\alpha^{i}(x^{i+1}-x^{i})$ (4)
+$=x+(x-1)\sum_{i=1}^{k}\alpha^{i}(x^{i})$ (5)
+$=x+(x-1)\frac{1-\alpha^{i+1}x^{i+1}}{1-\alpha x}.$ (6)
+C.2 Proof for Theorem 4.3
 Proof.
-Let $\alpha_{d_{1}},\alpha_{d2}$ and $c_{d_{1}},c_{d2}$ be the acceptence rate and cost ratio for model $d_{1},d_{2}$ respectively.
-Without loss of generality, we assume $\alpha_{d_{1}}>\alpha_{d2}$ and $c_{d_{1}}>c_{d2}$ Then the system satisfies $\alpha_{i}=f(i)\alpha_{d_{1}}+(1-f(i))\alpha_{d_{2}}$ and $c_{i}=f(i)\alpha_{d_{1}}+(1-f(i))c_{d_{2}}$ We prove using contradiction.
-Assuming all optimal systems need at least one interpolation of models at some positions.
-Let an optimal system of $f$ given $d_{1},d_{2}$, Without loss of generality, assume $r$ to be the first indices such that $f(r)$ not being an integer and the system $f$ to be the one with largest $r$ among all optimal systems.
-$T(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})=\frac{\alpha_{r}(1+\sum_{i=r+1}^{k}\prod_{j=r+1}^{i}\alpha_{j})+\sum_{i=0}^{r-1}\prod_{j=1}^{i}\alpha_{j}}{1+c_{r}+\sum_{i=1,i\neq r}^{k}c_{i}}$
-$=\frac{(f(r)\alpha_{d_{1}}+(1-f(r))\alpha_{d_{2}})(1+\sum_{i=r+1}^{k}\prod_{j=r+1}^{i}\alpha_{j})+\sum_{i=0}^{r-1}\prod_{j=1}^{i}\alpha_{j}}{1+f(r)c_{d_{1}}+(1-f(r))c_{d_{2}}+\sum_{i=1,i\neq r}^{k}c_{i}}$
-Let $A_{1}=1+\sum_{i=r+1}^{k}\prod_{j=r+1}^{i}\alpha_{j},A_{2}=\sum_{i=0}^{r-1}\prod_{j=1}^{i}\alpha_{j}$
-$=\frac{f(r)(\alpha_{d_{1}}A_{1}+A_{2})+(1-f(r))(\alpha_{d_{2}}A_{1}+A_{2})}{f(r)(1+c_{d_{1}}+\sum_{i=1,i\neq r}^{k}c_{i})+(1-f(r))(1+c_{d_{2}}+\sum_{i=1,i\neq r}^{k}c_{i})}$
-$\leq\max{(\frac{\alpha_{d_{1}}A_{1}+A_{2}}{1+c_{d_{1}}+\sum_{i=1,i\neq r}^{k}c_{i}},\frac{\alpha_{d_{2}}A_{1}+A_{2}}{1+c_{d_{2}}+\sum_{i=1,i\neq r}^{k}c_{i}})}$
-This means that $f(r)=0$ or $f(r)=1$ will result in a time improvement factor that is no less than that of a non-integer $f$, indicating that replacing $f(r)$ with an integer will still yield an optimal system.
-This contradicts the assumption that $f$ is the system with the largest $r$.
-As a result, there exists an optimal system that does not require interpolation.
-∎
-Figure 3: Acceptance rate of draft tokens by their positions on MMLU datasets, evaluated on FLAN-T5-small, base and large models.
+Let $\alpha^{\prime}=\alpha(\mathcal{M}_{d_{1}},\mathcal{M}_{d_{2}})$.
+We first calculate the expected number of tokens being generated in a step of vertical cascade with $\mathcal{M}_{d_{1}},\mathcal{M}_{d_{2}}$.
+With the property of generating function, the coefficient of term $x^{j}$ of $\phi^{n}(x)$ is the probability of the sum of acceptance length of $n$ speculative step being $j$.
+Therefore, $\phi^{n}(x)$ represents the probability generating function right before $\mathcal{M}_{t}$ performs the generation.
+To achieve the expected number of token acceptances of a probability-generating function, we seek for an operator that can map the probability-generating function into the desired expectation.
+To achieve the operator, we begin with a single polynomial term of $x^{j}$.
+Fortunately, given the end result of $\frac{1-\alpha^{j+1}}{(1-\alpha)}$ (Leviathan et al., 2023), the operator $T_{\alpha}(f(x))=\frac{1-\alpha f(\alpha)}{(1-\alpha)}$ will convert $x^{j}$ to $\frac{1-\alpha^{j+1}}{(1-\alpha)}$.
+In addition, due to the linearity of the operator, this can be extended to any polynomial.
+Therefore, we achieved the desired operator to map a probability-generating function into the desired expectation.
+Apply operator $T_{\alpha}$ to $\phi^{n}(x)$, we achieved the result of $\frac{1-\alpha\phi^{n}(\alpha)}{(1-\alpha)}$ for the expected number of accepted tokens.
+Furthermore, since the number of $\mathcal{M}_{d_{1}}$ calls is $n$, and $\mathcal{M}_{d_{2}}$ is called $k$ time for each $\mathcal{M}_{d_{1}}$ call given a total of $nk$ calls of $\mathcal{M}_{d_{2}}$.
+The time cost is $1+nc_{d_{1}}+nkc_{d_{2}}$ which implied the EWIF of the system being $\frac{1-\alpha\phi^{n}(\alpha)}{(1-\alpha)(1+nc_{d_{1}}+nkc_{d_{2}})}$.
+For Corollary 4.4, since both $10<\alpha<$ 1 and $10<\alpha^{\prime}<$ 1, we have $\alpha^{i+1}\alpha^{\prime i+1}<\alpha\alpha^{\prime}$, meaning that $\frac{1-\alpha^{k+1}\alpha^{\prime k+1}}{1-\alpha\alpha^{\prime}}<1$.
+Together with $\alpha-1<0$, we have $\phi_{\alpha^{\prime},k}(\alpha)<1+(\alpha-1)=\alpha$.
+If we also let $nkc_{d_{2}}=0$, we have $\frac{1-\alpha\phi^{n}(\alpha)}{(1-\alpha)(1+nc_{d_{1}}+nkc_{d_{2}})}>\frac{1-\alpha\alpha^{n}}{(1-\alpha)(1+nc_{d_{1}})}$, which is the EWIF for speculative decoding with step size $n$.

### v4

@@
 In cases where there is no match, we resort to a bigram model based on the probability distribution of Wikipedia (chosen to maintain the generality).
-Surprisingly, despite its simplicity, this straightforward and general algorithm has proven highly effective in aligning with neural language model outputs.
 We include a GPU-friendly version of the Max-Gram algorithm in Appendix A.
@@
 The algorithm requires an upper-triangular hyperparameter, $K_{nn}$, with each row serving as the stop criteria for a layer of recursive calls.
-For simplicity, we assume the lenience parameter $l$ is universal for the algorithm, except when the target model is under review; thus, the algorithm can benefit from the speedup of lenience without altering the output distribution.
+For simplicity, we assume the lenience $l$ is universal for the algorithm, except when the target model is under review; thus, the algorithm can benefit from the speedup of lenience without altering the output distribution.
 4 Analysis
@@
 6.2 Speculative Decoding
-Speculative decoding (Chen et al., 2023; Leviathan et al., 2023) is proposed to improve inference speed by using a draft model to perform auto-regressive decoding and a target model to check and replace the generated tokens.
-A key to further speed up is to increase the acceptance rate in speculative decoding, which can be done by aligning the distribution between the draft model and the target model.
-Some recent work has proposed to improve the performance of speculative decoding in a domain/task-specific setting.
+With the success of Speculative Decoding (Chen et al., 2023; Leviathan et al., 2023) in reducing the large language model inference latency, some recent works have attempted to improve Speculative Decoding by reducing the rejection rate.
 Zhou et al.
-(2023) use generalized knowledge distillation to finetune the draft model and achieve 10%-45% speedup compared to draft models without knowledge distillation.
-Liu et al.
-(2023) propose Online Speculative Decoding to continually update multiple draft models on observed user query data, and the draft models align with the query distribution in real-time.
-However, both approaches require additional task-specific training making it difficult to apply the methods in a general setting.
-On the other hand, He et al.
-(2023) propose constructing a datastore from a domain-specific corpus to serve as a draft model which can be used when a draft model is not available.
+(2023) propose using generalized knowledge distillation and achieve a lower rejection rate compared to other knowledge distillation methods.
+Avoiding an additional draft model, self-drafting is an approach to speculative decoding by reusing part of the target model together with added weight to perform drafting (Zhang et al., 2023; Hooper et al., 2024).
+Tree attention involves generating multiple candidates during drafting to increase the chance of acceptance (Spector & Re, 2023; Miao et al., 2024).
+Besides reducing the rejection rate, improving drafting efficiency can also reduce latency.
+Spector & Re (2023) propose using speculative decoding for drafting, showing similarities to the vertical cascade; however, their method only has two layers of speculative decoding and does not observe the recursive nature of the vertical cascade nor the lenience among draft models, two crucial aspects for the performance of vertical cascade.
 7 Conclusion

### v5

@@
 The Vertical Cascade eliminates autoregressive generation from neural models, while the Horizontal Cascade optimizes time allocation in drafting for improved efficiency.
-Combining both cascades, CS Drafting achieves up to an 81 percent additional speedup over speculative decoding in our experiments, while maintaining the same output distribution as the target model.
-Our code is publicly available at https://github.com/lfsszd/CS-Drafting.
-Machine Learning, ICML
-[…]
-[…]
+Combining both cascades, CS Drafting achieves greater speedup compared to the baselines in our experiments, while preserving the same output distribution as the target model.
+[^1] Code is publicly available at https://github.com/lfsszd/CS-Drafting.
 1 Introduction
-The advent of Large Language Models (LLMs), like GPT-4 (OpenAI, 2023), has marked a significant milestone in the field of natural language processing (NLP).
+The advent of Large Language Models (LLMs), like GPT-4 [17], has marked a significant milestone in the field of natural language processing (NLP).
 These models have not only excelled in various NLP tasks but have also found widespread applications in user-interactive settings, such as chatbots and virtual assistants.
@@
 Unfortunately, the ever-growing size of LLMs significantly increases the latency, especially in long-form generation, as autoregressive LLMs generate tokens one by one.
-An emerging solution, known as speculative decoding (Leviathan et al., 2023; Chen et al., 2023; Xia et al., 2023), shows potential to mitigate this issue.
+An emerging solution, known as speculative decoding [14, 4, 25], shows potential to mitigate this issue.
 In speculative decoding, a draft model (which is smaller and faster) generates $k$ tokens in each step (with $k$ being a hyperparameter) autoregressively, and these tokens are then reviewed by a target model (which is larger and slower) in parallel.
@@
 With a properly sized draft model, speculative decoding achieves a speedup of 2 to 3 times, making it a potential method for solving high latency issues.
-Figure 1: The figure shows an example of Cascade Speculative Drafting with target model $\mathcal{M}_{t}$ and draft models $\mathcal{M}_{d_{1}}$, $\mathcal{M}_{d_{2}}$, and $\mathcal{M}_{d_{3}}$.
-The horizontal cascade involves using larger draft models to generate the earlier tokens and smaller models to generate the later tokens.
-The vertical cascade requires each model to review drafts from smaller models with the exception of the smallest model which is a statistical language model.
-As the horizontal cascade and vertical cascade are orthogonal to each other, CS Drafting combines both approaches for optimal efficiency.
+Figure 1: The CS Drafting algorithm features a recursive and resource-efficient design, implemented through two cascades: the horizontal cascade and the vertical cascade.
+The horizontal cascade involves using larger draft models to generate the earlier tokens and smaller models for the later tokens.
+The vertical cascade requires each model to review drafts from smaller models with the exception of the smallest model, which is a statistical language model.
+As the horizontal cascade and vertical cascade are orthogonal, CS Drafting combines both approaches for optimal efficiency.
+The figure shows an example of Cascade Speculative Drafting with target model $\mathcal{M}_{t}$ and draft models $\mathcal{M}_{d_{1}}$, $\mathcal{M}_{d_{2}}$, and $\mathcal{M}_{d_{3}}$.
 However, since draft models are typically required to generate multiple tokens in multiple steps, where each generation still involves inefficient autoregressive decoding, the performance of speculative decoding could be limited by the drafting latency.
 This inefficiency is also indicated by Leviathan et al.
-(2023)’s experiments, where it was observed that very small models (e.g., around two orders of magnitude smaller than the target model) are usually the best choice for drafting because their inference cost is lower compared to that of a larger draft model, despite the fact that larger draft models usually have higher-quality generation.
+[14], where it was observed that very small models (e.g., around two orders of magnitude smaller than the target model) are usually the best choice for drafting because their inference cost is lower compared to that of a larger draft model, despite the fact that larger draft models usually have higher-quality generation.
 This underscores that improving drafting efficiency is crucial for further enhancing the performance of speculative decoding.
@@
 Nonetheless, we notice that it is not necessary to use only one draft model in speculative decoding—statistical language models can serve as the “draft” model for the neural draft model, thereby eliminating autoregressive generation from the neural draft model.
-Figure 2: The probability of acceptance of draft tokens in relation to their positions in a single step of speculative decoding, evaluated on FLAN-T5-small, base, and large models on GSM8K.
-The draft model generates 30 tokens at each step.
-The results on MMLU are shown in Figure 3 of the Appendix.
 Furthermore, our analysis in Figure 2 reveals a pattern during the drafting step: tokens generated later in the sequence by the draft model show a progressively lower probability of being accepted by the target model.
@@
 It indicates that later tokens from draft models are more prone to rejection, contributing less to the expected number of accepted tokens per draft step, yet incurring the same latency.
-Inspired by the above observations, we propose Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that comprises multiple draft models, starting with the smallest being a statistical language model.
+Figure 2: The probability of acceptance of draft tokens in relation to their positions in a single step of speculative decoding, evaluated on FLAN-T5-small, base, and large models on GSM8K and MMLU.
+The draft model generates 30 tokens at each step.
+Inspired by the above observations, we propose Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that comprises multiple draft models, with the smallest being a statistical language model.
 Each neural draft model reviews generations from a smaller model and then proposes its reviewed content to either a larger draft model or the target model.
-In this design, the drafting of each neural model will be accelerated by drafting from a smaller model, avoiding the inefficiency of autoregressive generation from neural models.
+In this design, the drafting of each neural model is accelerated by drafting from a smaller model, avoiding the inefficiency of autoregressive generation from neural models.
 We refer to this tiered speculative decoding approach as the Vertical Cascade.
@@
 2 Preliminary
-The core concept of speculative decoding (Leviathan et al., 2023) involves the utilization of a small draft model for sequential token generation with validation by a larger target model resulting in reduced latency.
+The core concept of speculative decoding [14] involves the utilization of a small draft model for sequential token generation with validation by a larger target model resulting in reduced latency.
 This design accelerates sampling from autoregressive models without altering output distributions.
@@
 At the end, the target model will generate one additional token following the accepted tokens.
-Such a design guarantees the output is the same as sampling autoregressively using the target model alone, as proven in Leviathan et al.
-(2023).
+Such a design guarantees the output is the same as sampling autoregressively using the target model alone [14].
 Speculative decoding was empirically validated on various tasks and model sizes, demonstrating a significant acceleration in inference times (2x-3x faster) compared to standard implementations, without affecting the outputs.
@@
 In this section, we introduce our proposed method, Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that incorporates two types of cascades: vertical cascade and horizontal cascade.
-Algorithm 1 CascadeSpeculativeDraftingStep
+{algorithm*}
+[!t] CascadeSpeculativeDraftingStep
   Require: draft models $\{\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{n}}\}$, target mode $\mathcal{M}_{t}$, $prefix$, flag $isFirstCall$, hyperparameters $K_{nn}$, $l$
@@
 We refer to this recursive speculative approach as Vertical Cascade.
-Additionally, we incorporate lenience, a hyperparameter that loosens the review process by the target model, allowing for faster speed at the trade-off of potentially differing results from the target model (Leviathan et al., 2023).
+Additionally, we incorporate lenience, a hyperparameter that loosens the review process by the target model, allowing for faster speed at the trade-off of potentially differing results from the target model [14].
 Lenience can be adopted during sampling or greedy decoding with speculative decoding.
@@
 [^2] For simplicity, we assume the probability outputs are not standardized.
-We refer the readers to Leviathan et al.
-(2023) for the discussion on standardized sampling.
+We refer the readers to the speculative decoding paper [14] for the discussion on standardized sampling.
 When performing greedy decoding, the acceptance condition becomes deterministic and is simply either $argmax\mathcal{M}_{d}(x)[i]=argmax\mathcal{M}_{t}(x)[i]$ or $\mathcal{M}_{d}(x)[i]\leq l\times\mathcal{M}_{t}(x)[i]$.
@@
 4 Analysis
-We begin with some notions. Let $\mathcal{M}_{t}$ be the target model, $\mathcal{M}_{d}$ be the draft model, and $k$ be the number of draft tokens generated per step.
+In this section, we provide theoretical analyses for Cascade Speculative Drafting.
+We begin with some notions.
+Let $\mathcal{M}_{t}$ be the target model, $\mathcal{M}_{d}$ be the draft model, and $k$ be the number of draft tokens generated per step.
 Expected acceptance rate $\alpha(\mathcal{M}_{t},\mathcal{M}_{d})$ is the probability of draft generation by $\mathcal{M}_{d}$ being accepted by target model $\mathcal{M}_{t}$.
@@
 Expected walltime improvement factor (EWIF) is the expected time improvement achieved by an algorithm under the i.i.d. assumption of token acceptance.
-Despite the simple setting of EWIF, Leviathan et al. (2023) have demonstrated that it aligns with the experimental results in most instances. Therefore, our analysis will concentrate on EWIF.
+Despite the simple setting of EWIF, it is demonstrated that it aligns with the experimental results in most instances[14]. Therefore, our analysis will concentrate on EWIF.
 4.1 Vertical Cascade
-We analyze EWIF of vertical cascade using generating functions, a well-studied topic in combinatorial mathematics (West, 2021).
-The properties of generating functions are useful in the recursion and evaluation process making the our final expression simple.
+We analyze EWIF of vertical cascade using generating functions, a well-studied topic in combinatorial mathematics [24].
+The properties of generating functions are useful in the recursion and evaluation process making our final expression simple.
 We begin with the derivation of the probability generating function for speculative decoding.
@@
 Corollary 4.4.
-$\phi_{\alpha^{\prime},k}(\alpha)<\alpha$ for any $01>\alpha>0,1>\alpha^{\prime}>0,k>$ 0, so if $c_{d_{2}}<<1$, the EWIF of $\mathcal{M}_{d_{1}}$ and $\mathcal{M}_{d_{2}}$ is higher than EWIF of $\mathcal{M}_{d_{1}}$ alone.
+$\phi_{\alpha^{\prime},k}(\alpha)<\alpha$ for any $01>\alpha>0,1>\alpha^{\prime}>0,k>$ 0, so if $c_{d_{2}}\ll 1$, the EWIF of $\mathcal{M}_{d_{1}}$ and $\mathcal{M}_{d_{2}}$ is higher than EWIF of $\mathcal{M}_{d_{1}}$ alone.
 Proof in Appendix C.2.
-Therefore, with the statistical model having negligible cost (i.e., $c_{d_{2}}<<1$), it can almost always improve the efficiency of an SD system.
+Therefore, with the statistical model having negligible cost (i.e., $c_{d_{2}}\ll 1$), it can almost always improve the efficiency of an SD system.
 4.2 Horizontal Cascade
@@
 Using the information provided by Leviathan et al.
-(2023), we calculate a simulated EWIF under the assumption that the event of acceptance by the target model is a Bernoulli trial.
+[14], we calculate a simulated EWIF under the assumption that the event of acceptance by the target model is a Bernoulli trial.
 The results, shown in Table 1, indicate that speculative sampling with a horizontal cascade achieved better EWIF than vanilla speculative sampling under this assumption.
@@
 BASE and SMALL refer to FLAN-T5-base and FLAN-T5-small.
-In the simulation, speculative sampling with horizontal cascade exceeded the performance of the vanilla speculative decoding on both CNN Dailymail (Nallapati et al., 2016) and WMT EnDe (Bojar et al., 2018) datasets.
+In the simulation, speculative sampling with horizontal cascade exceeded the performance of the vanilla speculative decoding on both CNN Dailymail [16] and WMT EnDe [2] datasets.
 5 Experiments
-5.1 Evaluation Metric
-Previous works on speculative decoding and related methods relied on walltime as an evaluation method.
-However, there are standardization and legitimacy concerns related to walltime.
-Leviathan et al.
-(2023) mentioned that cost coefficiency, the ratio between the time for a single run of draft model and target model will vary from system to system which can result in significant variation in wall time improvement across different systems, making comparisons across different research work difficult.
-In addition, a recent analysis suggests GPU speed can vary for the same GPU model with one being 1.5x faster than another GPU of the same model (Sinha et al., 2022); therefore, theoretical walltime improvement up to 1.5x can be undermined by GPU variation, and methods without any improvement can be measured with an improvement by running on a different GPU.
-The lack of standardization and stability makes walltime unfit for measuring algorithmic improvement.
-To resolve this issue, we design a new metric to measure speculative decoding and its related algorithms which can reduce the impact of varying/unstable GPU performance.
-Our proposed method, standardized walltime improvement (SWI), calculates the GPU times of the models, assuming that each run of a language model costs the same amount of time, an assumption made when inventing the speculative sampling algorithm (Chen et al., 2023).
-Given a speculative decoding system $S$ containing $k$ draft models $\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{k}}$, target model $\mathcal{M}_{t}$ solving a task $T$, and let $count(S,m,T)$ be the count of runs of model $m$ during the execution of task $T$ with system $S$, let $c_{i}$ be the expected cost coefficient being the time ratio between a single run of $\mathcal{M}_{i}$ and $\mathcal{M}_{t}$, and let $n$ be the total of tokens generated, the standardized walltime improvement is defined as
-$\begin{split}SWI&(\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{k}},\mathcal{M}_{t},T)=\\ &\frac{n}{count(S,\mathcal{M}_{t},T)+\sum_{i=1}^{n}count(S,\mathcal{M}_{d_{i}},T)c_{i}},\end{split}$
-which is simply the sum of GPU time by all models when $c_{i}$ is measured during a local experiment.
-The benefit of the $SWI$ evaluation is that $count(S,\mathcal{M}_{t},T)$ and $count(S,\mathcal{M}_{d_{i}},T)$ will be the consistent across different systems, so by using the same $c_{i}$ across different experiments the GPU performance variation will be eliminated; therefore, it provides a standardized, reproducible measurement.
-During our experiment, we will use two sets of $\{c_{i}\}$, existing $\{c_{i}\}$ reported by Leviathan et al.
-(2023) and an heuristic $\{c_{i}\}$ being the ratio of number of tunable parameters between $\mathcal{M}_{d_{i}}$ and $\mathcal{M}_{t}$.
-Dataset Algorithm $\{\mathcal{M}_{d_{i}}\}$ Speedup (MS) $k_{11}$ $k_{12}$ $k_{22}$ $l$
-GSM8K Autoregressive - 1 - - - -
-GSM8K S Decoding base $3.38$ 10 - - -
-GSM8K S Decoding small $3.06$ 11 - - -
-GSM8K CS Drafting base, mag $3.70$ 10 - - -
-GSM8K CS Drafting small, mag $3.19$ 11 - - -
-GSM8K CS Drafting base, small, mag 3.88 8 13 1 3
-MMLU Autoregressive - 1 - - - -
-MMLU S Decoding base $3.97$ 13 - - -
-MMLU S Decoding small $4.12$ 19 - - -
-MMLU CS Drafting base, mag $4.56$ 13 - - -
-MMLU CS Drafting small, mag $4.39$ 14 - - -
-MMLU CS Drafting base, small, mag 4.88 5 19 1 5
-Dataset Algorithm $\{\mathcal{M}_{d_{i}}\}$ Speedup (PW) $k_{11}$ $k_{12}$ $k_{22}$ $l$
-GSM8K Autoregressive - 1 - - - -
-GSM8K S Decoding base $2.99$ 8 - - -
-GSM8K S Decoding small $2.76$ 8 - - -
-GSM8K CS Drafting base, mag $3.27$ 9 - - -
-GSM8K CS Drafting small, mag $2.82$ 11 - - -
-GSM8K CS Drafting base, small, mag 3.43 5 9 1 3
-MMLU Autoregressive - 1 - - - -
-MMLU S Decoding base $3.42$ 10 - - -
-MMLU S Decoding small $3.51$ 11 - - -
-MMLU CS Drafting base, mag $4.21$ 6 - - -
-MMLU CS Drafting small, mag $3.99$ 13 - - -
-MMLU CS Drafting base, small, mag 4.32 5 8 1 5
+5.1 Experimental Setup
+Metrics We use both our proposed standardized walltime improvement and walltime for evaluation:
+  • Standardized walltime improvement (SWI) assumes each forward run of a model takes a constant amount of time which can be recorded data of previous work [14] or heuristics such as the total number of the parameters of a model.
+Under this assumption, the value of SWI is the speedup of the speculative method over autoregressive generation.
+SWI alleviates hardware variation and features full reproducibility of experiment results.
+  • Walltime refers to the actual elapsed time taken to complete a specific task or operation in a real-world scenario.
+Despite being less reproducible and sensitive to noise, walltime better represents the performance for individual users.
+In our experiment, walltime is measured in the form of the number of tokens generated per second on our GPU.
+Dataset Algorithm $\{\mathcal{M}_{d_{i}}\}$ Speedup (MS) Speedup (PW)
+GSM8K Autoregressive - 1 1
+[…]
+[…]
+[…]
+[…]
+[…]
+MMLU Autoregressive - 1 1
+[…]
+[…]
+[…]
+[…]
+[…]
 Table 2: The experimental results on FLAN-T5.
 Speedup (MS) is the standardized walltime improvement with the assumption that the latency of each run of a model is its number of parameters (model size).
-Speedup (PW) is the SWI with the assumption that the latency of each run of a model is the time cost data reported from previous work (Leviathan et al., 2023).
-$k_{11}$, $k_{12}$, $k_{22}$, $l$ are the hyperparameters.
-$k_{11}$ and $k_{12}$ represent the step limitation target model and the draft models, $k_{22}$ is the step limitations between the first and second draft model, and $l$ is lenience as shown in algorithm 1.
-For speculative decoding, the $k_{11}$ is simply the $k$.
-5.2 Setup
-To ensure the generality of our findings, we perform experiments on both encoder-decoder and decoder-only models.
-For encoder-decoder models, we choose our target and draft models from the FLAN-T5 (Chung et al., 2022) family for our experiment, as there is a large variation in model sizes within the FLAN-T5 family (ranging from 77 million to 11 billion parameters).
+Speedup (PW) is the SWI with the assumption that the latency of each run of a model is the time cost data reported from previous work [14].
+Datasets We chose two commonly used datasets for our experiments. For both datasets, we conducted experiments in a zero-shot chain-of-thought setup [13, 23]:
+  • GSM8K [7] is a dataset comprising 8,500 high-quality, linguistically diverse, grade-school math word problems.
+It focuses on multi-step reasoning with problems that are typically solvable using basic arithmetic in 2 to 8 steps.
+  • MMLU [10], or Massive Multitask Language Understanding, is a benchmark for testing how well large language models grasp knowledge.
+It encompasses 57 diverse subjects, ranging from elementary science to advanced law.
+Baselines To verify the effectiveness of both vertical and horizontal cascade strategies intuitively, we first compare the performance of CS Drafting with different numbers of cascades, as well as its performance against standard speculative decoding [14].
+Additionally, CS Drafting can also operate vertically by combining with other advanced decoding methods.
+To verify this, we also leverage tree attention, as used in Medusa [3], and compare its performance with Medusa.
+Implementation Details To ensure the generality of our findings, we perform experiments on both encoder-decoder and decoder-only models.
+For encoder-decoder models, we choose our target and draft models from the FLAN-T5 [6] family for our experiment, as there is a large variation in model sizes within the FLAN-T5 family (ranging from 77 million to 11 billion parameters).
 We use FLAN-T5-xxl as our target model, FLAN-T5-base and FLAN-T5-small as our reviewing draft models.
-For decoder-only models, we select LLAMA-2-chat-7B (Touvron et al., 2023) as the target model.
-Since there is no official release of a smaller model in LLAMA family, we use a pretrained 160M model with the same tokenizer as the reviewing draft model (Miao et al., 2024).
+For decoder-only models, we select Vicuna-7B [28], a fine-tuned version of LLaMA [20] as the target model.
+We use a 68M model with the same tokenizer as the reviewing draft model [^3] https://huggingface.co/double7/vicuna-68m.
+We also leverage tree attention [15, 3] with CS Drafting for the experiments on Vicuna-7B.
 In both cases, the Max-Gram algorithm is used as the generating draft model.
-To align our experiment with current common usage, we do not perform fine-tuning, and the generation is conducted in a zero-shot manner.
-Since we do not observe any significant difference between sampling with temperature $1$ and greedy decoding in previous speculative decoding experiments (Leviathan et al., 2023), and to ensure our experiments are fully reproducible, we perform sampling at temperature 00, i.e., using greedy decoding by default.
-Cascade Speculative Drafting Inference To reduce the number of hyperparameters to tune, we use MaG to generate 10 tokens at once, as it is rare for more than 10 tokens to be accepted.
-We do not use lenience when the reviewer is $\mathcal{M}_{t}$ to ensure the output distribution does not change.
-We also avoid lenience between MaG and its reviewer, since there is still a significant performance gap between MaG and a neural model.
-With these constraints, we are left with at most four hyperparameters: $k_{11}$, $k_{12}$, $k_{22}$, and $l$.
-For the CS Drafting step where the target model is the reviewer, $k_{11}$ and $k_{12}$ are used.
-$k_{21}$ and $l$ are used in the step where $\mathcal{M}_{d_{1}}$ is the reviewer.
-Baselines We consider both speculative decoding and vanilla autoregressive generation as our baselines.
-When performing speculative decoding with the target model as FLAN-T5-xxl, we opt for FLAN-T5-small and FLAN-T5-base as the draft models, as they perform better than FLAN-T5-large, according to Leviathan et al.
-(2023) and our preliminary experiments.
-When LLAMA-2-chat-7B is the target model, we use the 160M model as the draft model.
-We test a wide range of $k$-values, from 2 to 30, to ensure that we find the best result for each model serving as our draft model to establish our baseline.
-Dataset We chose two commonly used datasets for our experiments. For both datasets, we conducted experiments in a zero-shot chain-of-thought setup (Kojima et al., 2022; Wei et al., 2023):
-  • GSM8K (Cobbe et al., 2021) is a dataset comprising 8,500 high-quality, linguistically diverse, grade-school math word problems.
-It focuses on multi-step reasoning with problems that are typically solvable using basic arithmetic in 2-8 steps.
-  • MMLU (Hendrycks et al., 2021), or Massive Multitask Language Understanding, is a benchmark for testing how well large language models grasp knowledge.
-It encompasses 57 diverse subjects, ranging from elementary science to advanced law.
+Since we do not observe any significant difference between sampling with temperature $1$ and greedy decoding in previous speculative decoding experiments [14], and to ensure our experiments are fully reproducible, we perform sampling at temperature 00, i.e., using greedy decoding by default.
+To align our experiment with current common usage, we do not perform fine-tuning for CS Drafting, and the generation is conducted in a zero-shot manner.
+We include hyperparameter details in Appendix B.
+All of our experiments involving walltime are performed on a single NVIDIA A40 GPU.
 5.2 Experimental Results
-Comparison to Baselines Table 2 presents the experimental results.
-In two settings of SWI, Cascade Speculative Decoding has outperformed the speculative decoding algorithm.
-For GSM8K, Cascade Speculative Decoding achieved a maximum additional speedup of 44% over the fastest speculative algorithm; for MMLU, the maximum additional speedup improvement over speculative decoding is 81%.
-Overall, CS Decoding achieves a 3-4x speedup over autoregressive generation, establishing itself as a faster algorithm over the baselines.
+Table 2 presents the main experimental results.
+In two settings of SWI, Cascade Speculative Drafting has outperformed the speculative decoding algorithm.
+For GSM8K, CS Drafting achieved a maximum additional speedup of 44% over the fastest speculative algorithm; for MMLU, the maximum additional speedup improvement over speculative decoding is 81%.
 Effectiveness of MaG When comparing CS Drafting with one neural model and MaG against the fastest speculative decoding setup, we found that CS Drafting with one neural model gained up to a 70% speedup on MMLU and a 32% speedup on GSM8K.
-Notably, the MaG algorithm only involves a bigram model with as many parameters as the size of the tokenizer, making its memory cost negligible.
-Notably, the speedup gained using CS Drafting with one neural model involves no additional deployment overhead while reducing both latency and computational cost, making it a superior choice over speculative decoding.
+Notably, the MaG algorithm only involves a bigram model with parameters equal to the tokenizer size, making its memory cost negligible." In addition, the speedup gained using CS Drafting with one neural model involves no additional deployment overhead while reducing both latency and computational cost, making it a superior choice over speculative decoding.
 Draft Model Size Despite FLAN-T5-small mostly outperforming FLAN-T5-base as a draft model for speculative decoding, in CS Drafting with the aid of MaG, FLAN-T5-base consistently outperforms FLAN-T5-small.
 This implies that with the limitation of a single draft model, the ideal size of the draft model might increase with the assistance of the MaG model.
-Dataset Algorithm $\{\mathcal{M}_{d_{i}}\}$ Speedup $k_{11}$
-GSM8K Auto. - 1 -
-GSM8K S Decoding 160m $2.48$ 12
-GSM8K CS Drafting 160m, mag 2.86 15
-MMLU Auto. - 1 -
-MMLU S Decoding 160m $2.12$ 10
-MMLU CS Drafting 160m, mag 2.64 13
-Table 3: The experimental results on LLAMA.
-Results of Decoder-only Models As shown in Table 3, the results for decoder-only models exhibit similar traits to those observed in the experiments on encoder-decoder models.
-Therefore, for speculative decoding-based algorithms, we believe there is no distinct difference between an encoder-decoder model and a decoder-only model, aligning with the results in Leviathan et al.
-(2023).
+Algorithm GSM8K Walltime (Tokens/s) MMLU Walltime (Tokens/s)
+Autoregressive 33.34 33.11
+S Decoding 44.31 43.65
+CS Drafting 56.72 55.60
+Medusa 61.87 56.19
+CS Drafting + Tree Attention 63.81 63.37
+Table 3: The experimental results on Vicuna-7B.
+Results of Decoder-only Models As shown in Table 3, CS Drafting achieves a significant walltime improvement over speculative decoding.
+Moreover, CS Drafting exceeds Medusa when combined with tree attention [15, 3].
+This suggests that CS Drafting can be integrated with other efficient designs for speculative decoding to further accelerate inference.
+We leave the exploration of more advanced combinations as future work.
+Ablation Study We perform a hyperparameter study on $K_{00}$, the hyperparameter with the greatest effect on our experiments.
+As shown in Table 4, the performance of CS Drafting only decreases slightly when this hyperparameter is sub-optimal.
+Therefore, end users who do not require maximum performance can use a simple setup to achieve near-optimal performance with CS Drafting.
+Furthermore, we conduct an ablation study by removing the horizontal cascade.
+On GSM8K, with Vicuna-7B and $K_{00}=1$, this results in a performance drop from 56.16 to 53.55.
+$K_{00}$ Walltime (tokens/s)
+1 56.16
+2 55.51
+3 53.73
+Table 4: Results on GSM8K with Vicuna-7B under different generation length limits.
 6 Related Work
@@
 In the era of large language models, efficiency during inference becomes a key to model service.
-To reduce the model inference cost and speed up, several efficient methods have been proposed, including pruning, knowledge distillation and quantization (Treviso et al., 2023).
-Model pruning takes structured (Xia et al., 2022; Voita et al., 2019) or unstructured (Guo et al., 2021; Chen et al., 2020) methods to remove the redundant model parameters to reduce the storage memory and increase inference speed.
-Knowledge distillation takes the approach of transferring knowledge from a superior teacher model to a smaller student model (Hinton et al., 2015; Gou et al., 2021).
+To reduce the model inference cost and speed up, several efficient methods have been proposed, including pruning, knowledge distillation and quantization [21].
+Model pruning takes structured [26, 22] or unstructured [9, 5] methods to remove the redundant model parameters to reduce the storage memory and increase inference speed.
+Knowledge distillation takes the approach of transferring knowledge from a superior teacher model to a smaller student model [11, 8].
 Quantization maps high-precision data representations (e.g.
 32 bits) into low-precision ones (e.g.
-8 bits) to reduce memory consumption (Bhandare et al., 2019; Schaefer et al., 2023).
+8 bits) to reduce memory consumption [1, 18].
 6.2 Speculative Decoding
-With the success of Speculative Decoding (Chen et al., 2023; Leviathan et al., 2023) in reducing the large language model inference latency, some recent works have attempted to improve Speculative Decoding by reducing the rejection rate.
+With the success of Speculative Decoding [4, 14] in reducing the large language model inference latency, some recent works have attempted to improve Speculative Decoding by reducing the rejection rate.
 Zhou et al.
-(2023) propose using generalized knowledge distillation and achieve a lower rejection rate compared to other knowledge distillation methods.
-Avoiding an additional draft model, self-drafting is an approach to speculative decoding by reusing part of the target model together with added weight to perform drafting (Zhang et al., 2023; Hooper et al., 2024).
-Tree attention involves generating multiple candidates during drafting to increase the chance of acceptance (Spector & Re, 2023; Miao et al., 2024).
+[29] propose using generalized knowledge distillation and achieve a lower rejection rate compared to other knowledge distillation methods.
+Avoiding an additional draft model, self-drafting is an approach to speculative decoding by reusing part of the target model together with added weight to perform drafting [27, 12].
+Tree attention involves generating multiple candidates during drafting to increase the chance of acceptance [19, 15].
 Besides reducing the rejection rate, improving drafting efficiency can also reduce latency.
-Spector & Re (2023) propose using speculative decoding for drafting, showing similarities to the vertical cascade; however, their method only has two layers of speculative decoding and does not observe the recursive nature of the vertical cascade nor the lenience among draft models, two crucial aspects for the performance of vertical cascade.
+Spector et al.
+[19] propose using speculative decoding for drafting, showing similarities to the vertical cascade; however, their method only has two layers of speculative decoding and does not observe the recursive nature of the vertical cascade nor the lenience among draft models, two crucial aspects for the performance of vertical cascade.
 7 Conclusion
@@
 The vertical cascade eliminates the necessity of autoregressive generation from a neural language model, while the horizontal cascade effectively allocates the cost of drafting tokens at different positions.
-Our experiments show that CS Drafting achieves up to an 81 percent additional speedup over speculative decoding, while maintaining the same output distribution as the target model.
-Our work demonstrates that LLM inference can be further sped up by cascades without sacrificing generation quality or requiring additional task-specific training.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+CS Drafting achieves additional speedup over baselines in various settings while maintaining the same output distribution as the target model.
+8 Limitations
+Our experiments demonstrate strong performance for our methods.
+However, it is possible, though unlikely, that the outcome might differ on a system with different hardware configurations.
+To account for this, we report both standardized walltime improvement and raw walltime for a more robust evaluation.
+Additionally, we provide theoretical analyses to justify the improvements, which is independent of the hardware configurations.
+Acknowledgement
+This material is based upon work supported by the National Science Foundation IIS 16-19302 and IIS 16-33755, Zhejiang University ZJU Research 083650, IBM-Illinois Center for Cognitive Computing Systems Research (C3SR) and IBM-Illinois Discovery Accelerator Institute (IIDAI), grants from eBay and Microsoft Azure, UIUC OVCR CCIL Planning Grant 434S34, UIUC CSBS Small Grant 434C8U, and UIUC New Frontiers Initiative.
+Any opinions, findings, conclusions, or recommendations expressed in this publication are those of the author(s) and do not necessarily reflect the views of the funding agencies.
 Appendix A Max-Gram Implementation
@@
 26 return encoder_ids[:, index+1:index + n+1]
-Appendix B Additonal Results
-Figure 3: The probability of acceptance of draft tokens in relation to their positions in a single step of speculative decoding, evaluated on FLAN-T5-small, base, and large models on MMLU.
+Appendix B Hyperparameters
+To reduce the number of hyperparameters to tune, we use MaG to generate 10 tokens at once, as it is rare for more than 10 tokens to be accepted with the exception when CS Drafting is combined with tree attention.
+We do not use lenience when the reviewer is $\mathcal{M}_{t}$ to ensure the output distribution does not change.
+We also avoid lenience between MaG and its reviewer, since there is still a significant performance gap between MaG and a neural model.
+With these constraints, we are left with at most four hyperparameters: $k_{11}$, $k_{12}$, $k_{22}$, and $l$.
+For the CS Drafting step where the target model is the reviewer, $k_{11}$ and $k_{12}$ are used.
+$k_{21}$ and $l$ are used in the step where $\mathcal{M}_{d_{1}}$ is the reviewer.
+The results of experiments on encoder-decoder models with hyperparameters are shown in Table 5.
+When performing experiments with the decoder-only model, we fixed the hyperparameters of CS Drafting for different datasets to better align with most users who do not perform hyperparameter tuning.
+The k-matrix for CS Drafting is $[[2,10],[0,10]]$.
+When adding tree attention, we limit it to only the lead node with the highest probability of having children; the k-matrix is $[[1,3],[0,1]]$ with the number of children for each leading node being 8, while the other nodes have no children.
+Dataset Algorithm $\{\mathcal{M}_{d_{i}}\}$ Speedup (MS) $k_{11}$ $k_{12}$ $k_{22}$ $l$
+GSM8K Autoregressive - 1 - - - -
+[…]
+[…]
+[…]
+[…]
+[…]
+MMLU Autoregressive - 1 - - - -
+[…]
+[…]
+[…]
+[…]
+[…]
+Dataset Algorithm $\{\mathcal{M}_{d_{i}}\}$ Speedup (PW) $k_{11}$ $k_{12}$ $k_{22}$ $l$
+GSM8K Autoregressive - 1 - - - -
+[…]
+[…]
+[…]
+[…]
+[…]
+MMLU Autoregressive - 1 - - - -
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 5: The experimental results on FLAN-T5 with hyperparameter details.
+Speedup (MS) is the standardized walltime improvement with the assumption that the latency of each run of a model is its number of parameters (model size).
+Speedup (PW) is the SWI with the assumption that the latency of each run of a model is the time cost data reported from previous work [14].
+$k_{11}$, $k_{12}$, $k_{22}$, $l$ are the hyperparameters.
+$k_{11}$ and $k_{12}$ represent the step limitation target model and the draft models, $k_{22}$ is the step limitations between the first and second draft model, and $l$ is lenience as shown in algorithm 1.
+For speculative decoding, the $k_{11}$ is simply the $k$.
 Appendix C Proof
@@
 To achieve the operator, we begin with a single polynomial term of $x^{j}$.
-Fortunately, given the end result of $\frac{1-\alpha^{j+1}}{(1-\alpha)}$ (Leviathan et al., 2023), the operator $T_{\alpha}(f(x))=\frac{1-\alpha f(\alpha)}{(1-\alpha)}$ will convert $x^{j}$ to $\frac{1-\alpha^{j+1}}{(1-\alpha)}$.
+Fortunately, given the end result of $\frac{1-\alpha^{j+1}}{(1-\alpha)}$ [14], the operator $T_{\alpha}(f(x))=\frac{1-\alpha f(\alpha)}{(1-\alpha)}$ will convert $x^{j}$ to $\frac{1-\alpha^{j+1}}{(1-\alpha)}$.
 In addition, due to the linearity of the operator, this can be extended to any polynomial.